### v1

# Perception of Visual Content: Differences Between Humans and Foundation Models

Abstract
Human-annotated content is often used to train machine learning (ML) models. However, recently, language and multi-modal foundational models have been used to replace and scale-up human annotator’s efforts. This study compares human-generated and ML-generated annotations of images representing diverse socio-economic contexts. We aim to understand differences in perception and identify potential biases in content interpretation. Our dataset comprises images of people from various geographical regions and income levels washing their hands. We compare human and ML-generated annotations semantically and evaluate their impact on predictive models. Our results show low similarity between human and machine annotations from a low-level perspective, i.e., types of words that appear and sentence structures, but are alike in how similar or dissimilar they perceive images across different regions. Additionally, human annotations resulted in best overall and most balanced region classification performance on the class level, while ML Objects and ML Captions performed best for income regression. Humans and machines’ similarity in their lack of bias when perceiving images highlights how they are more alike than what was initially perceived. The superior and fairer performance of using human annotations for region classification and machine annotations for income regression show how important the quality of the images and the discriminative features in the annotations are.
Introduction
Machine Learning (ML) is prevalent in numerous applications, e.g., facial recognition and recommendation systems. However, many of these use cases could potentially result in discrimination and unfair decisions. For example, it is possible that a facial recognition algorithm is less likely to perform well when predicting people of certain ethnic groups, e.g., people of African heritage may not be identified as accurately as people of European origin. This is problematic, especially as we aim to rely more on ML for daily tasks. Since supervised ML models require data for training, biased training data could affect the fairness and bias of the learned model – also known as algorithmic bias – which can cause it to generate prejudiced predictions. Unfortunately, it is quite likely that the data will contain bias considering that they are commonly annotated by human annotators (Demartini, Roitero, and Mizzaro 2024).
Furthermore, generative models have started to be used to produce labelled data for training and evaluation of ML models, which amplifies this issue (Thomas et al. 2023). It is also unclear how these machine annotations are different to human annotations. Since human annotation could contain biases, one possible approach to creating potentially less biased annotations is to use machine-generated annotations. These annotations could likely be more objective compared to those generated by humans.
In this study, we first investigate how similar annotations generated by humans and machines are, and explore possible justifications for their similarities or differences, in the context of how they think and perceive images. This leads to our first research question (RQ1): “How similar are annotations generated by humans compared to the annotations generated by ML models?” Moreover, due to the aforementioned biases that could appear in the annotations, we look into how different annotations affect fairness in predictive models. This leads to our second research question (RQ2): “How do different combinations of annotations affect fairness in ML predictive models?” To do this, we use the annotations in vector representation to train predictive ML models. These annotations are based on images of people washing their hands from varying geographical regions and socio-economic classes.
This study contributes to the fields of information retrieval and algorithmic bias in several ways. Firstly, it provides a comparative analysis of human-generated and machine-generated annotations in the context of socio-economic diverse content, which give insights on how humans and machines are alike in their bias when perceiving visual content despite their low-level differences, filling a gap in the existing literature. Secondly, it explores whether different annotations sets can impact the efficacy and fairness of ensemble-based ML models, thus advancing both theoretical and practical understanding of this critical issue.
Related Work
Effect of Bias on ML Models. It is inevitable that human annotations for training ML models may contain bias and stereotypes. Numerous qualitative and quantitative studies have demonstrated that there exist biases in numerous techniques, datasets, and models related to Natural Language Processing and Information Retrieval (Baeza-Yates 2020; Bolukbasi et al. 2016; Caliskan, Bryson, and Narayanan 2017). These biases could reinforce stereotypes that already exist in society (Rekabsaz, Kopeinik, and Schedl 2021). Fan et al. (2022) investigated this by conducting a study in which participants were asked to annotate video content associated with various regions and levels of income. When sentiment analysis was performed on the annotations, it was found that they contained regional and income-level biases. Unfortunately, as discussed in their paper, one of the main concerns of these biased annotations is that they impact ML models when used as training data. This causes various search and recommendation systems to use these algorithms to feed users with stereotypical content. Additionally, crowd annotators may be experts in different domains, so high-quality annotations are not always guaranteed. Not only could noise be introduced, but the annotated data may not be consistent (Cao et al. 2023). These unreliable annotations may also introduce more biases, which would affect the ML model’s performance. Devillers, Vidrascu, and Lamel (2005) emphasise how these biases can be dangerous when the training dataset involves actual humans as these learned biases tend to mimic forms of discrimination common amongst humans, such as racial and gender bias. An example of this is Microsoft’s “Tay”, an AI chatbot that initially mimicked a teenage girl but became anti-semitic, racist, and sexist over a span of only 16 hours (Zemčík 2020). Sun, Nasraoui, and Shafto (2020) further discuss how the problem worsens when users make decisions based on the biased content provided by these models. These decisions could then result in data used to train other algorithms in the future, thereby creating a never-ending bias-reinforcement cycle. In order to mitigate this issue and break the loop, more diverse data is needed (Fan et al. 2022). Chhikara et al. (2023); Dash et al. (2019) further support this by discussing how textual content written by numerous different authors will contain a greater variation of perspectives, especially on polarising topics, such as religion and politics. This makes it more crucial to see how fairly different perspectives are reflected when attempting to summarise these into one text. van Atteveldt, van der Velden, and Boukes (2021) looked at comparing sentiment analysis methods in news headlines. They used various approaches to compare the sentiment scores produced for these headlines, including manual annotation, crowd-coding, dictionary approaches, and ML algorithms. In this case, their ML algorithm was used to evaluate the news headlines rather than create equivalent annotations to compare against the human annotations. As compared to this body of literature, we focus on comparing human and machine-generated annotations of visual content with the goal of understanding how bias may differ between these annotations.
Dealing with Imbalanced Data. It is important for data to be well-balanced, ensuring that every class in the dataset is well-represented. This is especially true for traditional ML models as they were originally designed for datasets that are balanced. There are various ML methods that were designed to deal with imbalanced datasets, such as random under-sampling and over-sampling techniques (Estabrooks, Jo, and Japkowicz 2004; Chawla et al. 2002), creating synthetic data (Chen et al. 2023; He et al. 2008; Han, Wang, and Mao 2005), and cost-sensitive learning where the model is penalised more when a minority class sample is incorrectly classified compared to a majority class sample (Ting 2002). In our work, we adopt under-sampling techniques to overcome the issue of data imbalance across some classes in our dataset due to the limited number of available samples.
Ensemble Models. Minimising the trade-off between accuracy and fairness is a question that is very relevant to our work. There has been research indicating that a loss in accuracy is inevitable when a model is trained with limitations relating to fairness. Based on Kenfack et al. (2021), we know that compared to individual classifiers, ensemble models have an improved predictive performance, but compromise on fairness. In our work, we include the use of ensemble models in our experiment to test whether there is the same effect on our dataset.
Cognitive Differences in Humans and Machines. To this day, human experts are still needed to provide accurate and reliable annotations that could be used to train ML models (Valizadegan, Nguyen, and Hauskrecht 2013). Understanding how differently humans and machines think is crucial in helping us develop better and fairer automated systems that could assist us in solving problems, i.e., annotating images for ML. Sharifi Noorian et al. (2022) proposed Scalpel-HS, a framework that is intended to identify unknown unknowns in images by combining human intelligence and machine learning. Unknown unknowns refer to instances that an ML model fails to identify or has not been trained on. This framework aims to reduce cognitive load for humans by using ML models to first identify relevant objects and relationships, so that the humans can focus on evaluating relevance rather than performing the task from scratch. In our study, instead of using ML to aid humans in the annotating process, we investigate how similar human annotations and machine-generated annotations are and factors that explain their similarity or dissimilarity considering how humans and machines perceive images.
How machines think inevitably affects their decision-making. In the context of image recognition tasks, Zhang et al. (2021) investigates how humans and ML models select features differently. Their study revealed that humans are not always better than ML models at selecting discriminative features for image recognition and that humans rely on contextual information, which may affect how they select the features. In our work, instead of image segmentation tasks, we look at how humans and machines perform annotation tasks on images. From this, we explore how well the discriminative features in the annotation sets are by using them to train predictive models.
Methodology
This section details the methodology we used to address the research questions introduced earlier. The methodology begins with data collection from Dollar Street. Annotations are preprocessed and added to the data either through crowdsourcing or ML techniques. The methodology is divided into two key experiments: the first evaluates annotation similarity; the second investigates the impact of different annotation sets on supervised models, using ensemble learning and comparing the performance of human and machine annotations.
Experimental Set-Up
In this study, we utilised CPU resources to conduct all experiments. The computations were performed locally on a device with an Apple M3 Pro Chip, 11 CPU cores, and 36GB of RAM.
Dataset
The dataset utilised in our study was obtained from Dollar Street, a project by Gapminder, and is available under a Creative Commons license CC BY 4.0 (Gapminder 2021). Dollar Street visualises global public data by arranging families based on their income levels, from the poorest on the left to the wealthiest on the right. It features photographs of everyday objects from more than 240 families in 50 countries, providing a representation of living conditions across different economic classes. There are 197 unique instances of images and videos utilised in our experiments, consisting of people washing their hands, sourced from various geographical regions and income levels. The topic of hand washing was selected due to its crucial role in public health and its practice across all cultures and income levels.
Data Annotation
Figure 1: Handwashing images annotated using the pre-trained Faster R-CNN model.
Crowdsourcing. The first annotation set is sourced from a structured crowdsourcing process conducted on Amazon Mechanical Turk (MTurk) by Fan et al. (2022). In their study, crowd workers viewed short videos of individuals washing their hands from various global regions and income levels, providing textual annotations. These annotations captured various aspects of handwashing behaviour and the surrounding environmental context. The dataset includes annotations for 112 videos, each annotated by 10 US-based workers. Fan et al. (2022) implemented quality control measures to ensure reliable annotations. Consent from the participants was obtained through a signed consent form. Additionally, no personally identifiable information or offensive content were collected in the crowdsourcing process. We use these existing human annotations in our current study, benefiting from their comprehensive approach to data collection.
ML Object Detection. Images used in this study are retrieved frame-by-frame from videos capturing handwashing practices. The images are processed using a pre-trained Faster R-CNN model with a ResNet-50-FPN backbone, as outlined by Li et al. (2021), which offers improvements over the original model configuration introduced in Ren et al. (2017). This model is available under the MIT License. This model was originally trained on the COCO dataset (Lin et al. 2014) and is utilised to generate our machine-based object annotations, as illustrated in Figure 1. We selected the Faster R-CNN model for its prominence in object detection, its efficiency, and its effectiveness (i.e., mean average precision) on the COCO labels  (Mohammed, Tannouche, and Ounejjar 2022; Ren et al. 2017). We chose to perform object detection on the dataset as most of the images contain objects, i.e., bottles, sinks, instead of subjects performing actions, i.e., person washing hands. Due to this, the COCO dataset was chosen as the most suitable for training as it contains annotations of everyday objects.
ML Captioning. The images are also passed through a pre-trained BLIP model (Li et al. 2022), with Vision Transformer (ViT) base backbone (Dosovitskiy et al. 2020), for image captioning purposes. The model is available under the BSD 3-Clause ”New” or ”Revised” License. The model was originally trained on the COCO captions dataset (Chen et al. 2015), which helps ensure that this and the object detection model are trained on data that are based on the same images. The BLIP model configuration was chosen as it has achieved state-of-the-art results, outperforming other Vision-Language Pre-training (VLP) methods. We chose to perform image captioning as it would likely resemble the annotations generated by humans.
Data Preparation and Preprocessing
The data cleaning process constituted the initial step in preparing our dataset for analysis. In this phase, corrupted images were identified and removed based on criteria including integrity checks and visual inspections. Images that did not contain labels from any of the annotation categories are also removed, i.e., images that did not have any objects detected in it. This process reduced the original dataset to 197 unique data samples with distinct IDs.
The annotations were preprocessed to deal with inconsistencies in the human annotations. These steps involve noise/punctuation removal, conversion to lower case, auto-correcting misspellings, and segmenting strings that are concatenated.
We utilised a sentence embedding model to convert the three types of annotations into sentence embeddings. Instead of using a pre-trained sentence transformer model to generate the embeddings, we trained a sentence transformer model from scratch on all three sets of annotations. We did this as we want the model and representations to be very focused, and due to the limited number of samples, fine-tuning an existing sentence transformer, such as sBERT, resulted in worse performance.
To create the sentence transformer, we used a pre-trained Transformer model, “distilroberta-base”, which is available at the Hugging Face Hub (Sanh et al. 2019), to obtain the word embedding model. We then combined this word embedding model and a mean pooling function over the token embeddings so that it could generate a fixed-size vector as output. Using the sentence transformer model, we generated embeddings with a size of 384 dimensions.
Experiment 1: Investigating Annotation Similarity (RQ1)
To investigate the similarity between human, machine-generated object labels, and machine-generated captions, we utilised the embeddings that have been generated for the three types of annotations across all images.
In the first part of this experiment, we utilised T-Stochastic Neighbor Embedding (T-SNE) to visualise the vector embeddings in two-dimensional space (van der Maaten and Hinton 2008). Doing this helps us analyse the distances of the vectors, and thus the annotations, for every image in the dataset and see how they are distributed.
We then used the sentence transformer model’s built-in similarity function to compute the cosine similarity scores between each possible pair of annotations. Using these scores, we look at the average similarity scores for each pair and see how the pairwise scores are distributed using scatterplots and boxplots.
These experiments are appropriate for addressing RQ1 as t-SNE and computation of similarity scores across pairs allows us to evaluate how similar the annotations are, whereas the pairwise scatter plot helps us derive trends and patterns between annotations, further evaluating their likeness to one another.
Experiment 2: Exploring How Annotations Impact Fairness in Predictive Models (RQ2)
To investigate how the three annotation sets affect predictive models, we use the vector embeddings of every annotation as training data for two supervised learning models.
Before training the models, the dataset was split into training and test sets with a split ratio of 80:20, respectively . The training set was first used to run Grid Search Cross-Validation (CV) with each of the two models to obtain the optimal parameters for each model whilst performing 5-fold cross-validation. Grid Search CV was performed in every independent run.
The first model is an RUS Boost Classifier used to classify each sample on region (Seiffert et al. 2008). This classifier involves the incorporation of the Adaptive Boost algorithm, which is a sequential ensemble technique that can lower a model’s bias and variance, with random under-sampling, which helps deal with data imbalance (Mohammed and Kora 2023; Seiffert et al. 2008). This makes the algorithm ideal for our dataset due to its small size and there not being an equal number of samples from every region. The base estimator used in the model is the Decision Tree Classifier (Swain and Hauska 1977). After obtaining the optimal parameters, i.e., maximum number of estimators where boosting stops and the learning rate, for the model using grid search CV, we trained the classifier using these parameter values. The trained classifier is run 10 times with different random seeds to obtain a mean F1 score for each class (region) and a mean overall weighted F1 score, with their respective standard deviations. This is done to evaluate the model effectiveness for each class and the overall model performance. The model is run multiple times to allow us to evaluate its stability. The Welch’s t-test was performed to compare whether the difference in performance of the best and worst overall-performing models are statistically significant. This test was performed due to the Gaussian distribution of the F1-Scores and varying variances between the models.
The second model uses the Adaboost.R2 algorithm to predict each sample’s income (Drucker 1997). Similar to the region classifier, this algorithm uses a Decision Tree Regressor as its base estimator. Since boosting helps reduce prediction error on the small dataset, this algorithm is suitable for this task. Using grid search CV, we optimised for the number of estimators, learning rate, and the type of loss function used in the Adaboost.R2 algorithm. The model was executed 10 times with different random seeds. The root mean square error (RMSE) was obtained for every run, both for overall and region-based evaluation. The RMSE takes the square root of the mean square error, which performs the measurement using the same units as the target variable, i.e., income level. A regression plot is generated to visualise the distribution of predicted samples by income for the best-performing model. The Kolmogorov-Smirnov test was performed to compare whether the difference in performance of the best and worst overall-performing regression models are statistically significant. This test was performed due to the non-Gaussian distribution of the RMSE scores between the models.
These predictive models are appropriate in addressing RQ2 as it helps us investigate how often each region is correctly identified for a given image, and how well income is predicted for images in a given region. For evaluation purposes, the F1-score is appropriate and sufficient as it is a good metric for evaluating classification performance on imbalanced datasets, as ours is, whereas the RMSE scores are appropriate and sufficient as the measurement being done in the same units as the target variable makes it easier to interpret. Looking at these evaluation metrics for every region helps us determine whether the models perform better for particular regions, giving insight on how fair the model is.
Results and Discussion
RQ1: How similar are human-generated and ML-generated annotations?
Figure 2: t-SNE Visualisation of Sentence Embeddings of Each Label
Annotations Used Mean ± s.d.
ML Objects and ML Captions 0.87 ± 0.05
ML Objects and Human Labels 0.42 ± 0.1
ML Captions and Human Labels 0.45 ± 0.06
Table 1: Mean and standard deviation of similarity scores between each pair of annotations across all samples.
Figure 3: A plot of pairwise relationships between each pair’s similarity score.
Figure 2 illustrates a t-SNE visualisation of the vector embeddings of the three annotation categories for each data sample. The results show how, despite having trained an embedding model from scratch, the embeddings still form into three distinct clusters. Further analysis shows that the human labels cluster is furthest from the ML captions cluster, with the ML object labels cluster in between them.
The clusters of embeddings are due to differences in the annotations. Human annotations are holistic, encapsulating both the objects and the array of actions unfolding in the captured images and videos. On the other hand, the ML-generated object labels are primarily object-centric. The COCO dataset, which the ML object detection model is trained on, does not extend to action recognition, which is not relevant due to the object-oriented images in the dataset, thereby inherently limiting the ML-generated object annotations to object identification. The ML-generated captions, although more descriptive than the object labels, tend to not be as specific as the human annotations since they are more likely to contain easily recognisable actions and objects rather than complex descriptions of people’s emotions or the environment. Additionally, there are apparent syntactic differences between the annotations - the ML object labels consists of lists of objects per image; the ML captions consist of a phrase per image; while the human labels are a list of words and phrases. It is also important to note that a limitation that could have affected this distinct categorisation is that stop word removal and lemmatisation were not performed during preprocessing as to preserve nuances and context that contain relevant information. This causes the sentence structures in the annotations to remain unchanged, resulting in more pronounced categorisations. Considering these differences, the distinct clusters in the visualisation are expected.
However, since t-SNE is only a two-dimensional representation of the data, it would be useful to also look at the similarity scores, generated by the BLIP model’s cosine similarity function, between each pair of annotations across the dataset. This is shown in Table 1. These values give a more accurate and quantitative comparison of the annotation similarities. It can be seen that the ML object labels and the ML captions are the most similar pair of annotations (0.87), followed by ML captions and human labels (0.45), and ML object labels and human labels (0.42). It is important to note that when comparing human labels with ML-generated ones, the similarity scores are relatively low, regardless of the type of ML annotation.
The high similarity of both ML annotation types is a reasonable outcome. Although the annotations were generated by two different pre-trained ML models which were trained with different types of training data, i.e., labels versus image-text pairs, the training data both originated from the COCO dataset and are based on the same set of images. This means that there would be more overlap of objects between the two annotation categories. In contrast, the low similarity values between either ML annotations with the human labels is expected since there may be higher variety of words and phrases that were used by the human annotators. Moreover, when we compare the t-SNE visualisation and the similarity scores, it may seem inconsistent at first glance due to the ML objects and human labels pair having the lowest similarity score despite being seemingly furthest apart visually. However, since t-SNE is a dimensionality reduction technique, there may be features that were lost in the process, resulting in less accurate representations.
To further explore the relationships of these similarity scores, a plot of pairwise relationships is presented in figure 3. The plot illustrates how similarity scores between two annotation categories correlate to another pair’s scores for a given image. The plot is colour-coded by region to allow further analysis of similarity scores by geographical location. The analysis reveals that there are no particular trends or patterns with respect to region. Moreover, when looking at the subplots for ML objects-ML captions pairs, i.e., subplots on first column/row, the similarity scores tend to be relatively high, on average, with the exception of several outliers. This aligns with the previous observation that ML objects and ML captions have the highest average similarity score (0.87). This pair also has the smallest variance out of all pairs, which is also evident when comparing the vertical spread of scores across the subplots in all rows, i.e., the scores of the ML objects-ML captions pair are not spread as much vertically when disregarding the outliers. In contrast, the ML objects-human labels pair, i.e., second row/column, appears to have the highest variance, which also aligns with the standard deviation of the mean similarity score observed in the previous experiment (0.1). It is also shown that the ML objects-ML captions pair has the same outliers as the ML objects-human labels pair (first row, second column), but not with the ML captions-human labels pair (first row, third column). Interestingly, the ML captions-human labels pair (third column/row) seems to have no outliers.
The lack of discernible patterns in the way the scores are distributed across regions suggests that the images have little to no effect on how similar two sets of annotations will be. This could mean that the discriminative features in the images, or lack thereof, are detected equally by both humans and ML models. Therefore, none of the annotation sets describe images from a particular region any better or worse than the other. This also suggests that humans and machines may be similar in terms of how biased they are when perceiving images.
The low variance of similarity scores for the ML objects-ML captions pair suggests that for the majority of images, both annotations consistently contain similar objects, which - similar to the previous discussion on their high similarity scores - is expected due to the origin of the both models’ training data. Similarly, the high variance of scores for the ML objects-human labels pair aligns with expectations. This is due to, not only syntactic differences, i.e., objects vs longer phrases or sentences, but also differences in the way machines and humans interpret visual content, i.e., machines are more likely to capture easily recognisable objects and actions but may fail to capture more complex and nuanced observations, such as environment aesthetics or partially obstructed objects.
Moreover, the fact that the ML objects-ML captions pair has the same outliers as the ML objects-human labels pair only (first row, second column) indicates that for those particular images, the ML object detection model may be detecting objects that do not appear in the corresponding ML captions or human labels, or failing to identify objects that do. This reasoning is consistent with the findings of Zhang et al. (2021), which suggest that deeper neural networks, such as ResNet - the architecture of our object detection model - may struggle more with difficult images. Perhaps there is obstruction or lack of objects in the image, leading to no or incorrect objects being detected in these outliers, resulting in the low similarity scores when compared to their respective ML Captions and human labels. The ML captioning model as well as human annotators have the advantage in that they are able to capture non-object elements, such as actions, e.g., “wash”, or environmental descriptions, e.g., “dirty”.
RQ2: How do different combinations of annotations affect fairness in ML predictive models?
F1-Score (Mean ± Standard Deviation)
Annotations Used Africa Asia Europe The Americas Overall
ML Objects 0.43 ± 0.05 0.45 ± 0.06 0.30 ± 0.09 0.51 ± 0.10 0.42 ± 0.03
ML Captions 0.60 ± 0.05 0.32 ± 0.05 0.57 ± 0.02 0.38 ± 0.06 0.45 ± 0.02
Human Labels 0.62 ± 0.05 0.54 ± 0.13 0.57 ± 0.05 0.51 ± 0.11 0.56 ± 0.07
ML Objects & ML Captions 0.64 ± 0.07 0.50 ± 0.05 0.49 ± 0.06 0.46 ± 0.09 0.53 ± 0.04
ML Objects & Human Labels 0.57 ± 0.05 0.41 ± 0.05 0.57 ± 0.06 0.55 ± 0.10 0.51 ± 0.04
ML Captions & Human Labels 0.60 ± 0.09 0.38 ± 0.05 0.62 ± 0.07 0.51 ± 0.12 0.51 ± 0.03
ML Objects & ML Captions & Human Labels 0.65 ± 0.06 0.46 ± 0.04 0.51 ± 0.06 0.53 ± 0.11 0.53 ± 0.05
Table 2: Comparison of the region classification model performance across different annotation types. The mean and standard deviation of the F1-Score were recorded over 10 independent runs, both for the overall model performance and for each class individually.
RMSE (Mean ± Standard Deviation)
Annotations Used Africa Asia Europe The Americas Overall
ML Obj 256.17 ± 88.2 1161.35 ± 116.23 2467.03 ± 250.75 1142.27 ± 83.42 1341.65 ± 85.56
ML Captions 179.69 ± 54.86 1456.47 ± 132.90 2164.40 ± 99.97 1815.89 ± 59.91 1564.56 ± 48.54
Human 228.35 ± 55.14 1657.22 ± 107.99 2675.59 ± 104.75 1433.51 ± 74.91 1633.25 ± 49.57
ML Obj & ML Captions 168.79 ± 39.57 1223.70 ± 137.91 2111.71 ± 138.83 1191.59 ± 136.96 1279.39 ± 57.66
ML Objects & Human 117.97 ± 11.04 1497.18 ± 86.39 2621.28 ± 219.83 1108.55 ± 85.18 1480.72 ± 78.63
ML Captions & Human 149.31 ± 16.17 1523.02 ± 108.62 2525.53 ± 186.66 1202.37 ± 72.59 1486.67 ± 86.93
Obj & Captions & Human 125.85 ± 8.27 1407.31 ± 54.97 2549.58 ± 135.91 1191.46 ± 82.55 1449.52 ± 45.33
Table 3: Comparison of the income regression model performance across different annotation types. The mean and standard deviation of the RMSE were recorded over 10 independent runs, both for the overall model performance and for each class individually.
This subsection aims to compare the efficacy of various annotation combinations - ML Objects, ML Captions, and Human Labels - to enhance the performance of models devised for geographical region classification and income regression. The performance of the models was appraised using the F1 score, which balances precision and recall, and Root Mean Square Error (RMSE), which measures performance using the same units as the target variable.
Table 2 tabulates the F1-Scores of the region classification ensemble model across various combinations of the three annotation sets: ML Objects, ML Captions, and Human Labels. The mean and standard deviation of the F1-Scores across 10 independent runs were recorded, both for the overall model and for each class (region).
It is discerned that the Human Labels resulted in the highest overall model performance (0.56). The models utilising combinations of the human and ML annotations exhibited poorer overall performance, followed by the ML Captions Model (0.45) and the ML Objects Model (0.42). Compared to other models, the Human Labels model has the smallest range of F1-Scores across all classes, ranging from 0.51 to 0.62. Moreover, when looking at per-class performance, it is revealed that each class best performs with different models. Africa (class) exhibits the highest performance with the model that uses all three annotation sets (0.65), Asia with the Human Labels Model (0.54), Europe with the ML Captions and Human Labels Model (0.62), and the Americas with the ML Objects and Human Labels Model (0.55). One notable observation is that regardless of overall or per-class performance, human labels are always used in the models - none of the ML-only models resulted in superior overall or per-class performance.
The superior overall performance of the Human Labels model is expected. As discussed by Zhang et al. (2021), humans are more likely to use contextual reasoning in their decision-making process. Although the paper talks about this in the context of humans segmenting images, this may similarly extend to annotation tasks, positively impacting the model’s performance in classifying regions. Perhaps it is the case that classifying regions in particular requires more than descriptions of objects and actions, both of which are what the ML models were trained on. This could likewise influence the fairer and more balanced class-level performance of the Human Labels model, as evidenced by the range of the F1-Scores across all classes. This may also account for why the best-performing models, both overall and class-level, are those that involve human labels.
While Zhang et al. (2021) claim that humans are not necessarily better than machines in choosing discriminative features for image segmentation, our classification model results highlight a different perspective. It is evident that the usage of human labels, whether on its own or in combination with another annotation set, leads to more accurate region classifications. From this, it can be deduced that humans are, in fact, better than machines at creating image annotations that contain these discriminative features in the context of region classification.
Using the Welch’s t-test, we found that there are statistically significant differences between the overall F1-scores from 10 independent runs across the Human Labels model and the ML Objects model at $p$-value $<$ 0.001.
Table 3 shows the Root Mean Square Error (RMSE) for the AdaBoost Regression algorithm used to predict income. Similar to the classification model, the experiment was conducted over 10 independent runs across different combinations of the three annotation sets. The mean and standard deviation of the RMSE values were obtained for the overall model and for each region.
It is apparent that, unlike the classification model, the regression model with ML Objects and ML Captions performed the best overall, indicated by its lowest RMSE value (1279.39). This model also has the smallest range of RMSE values across all classes, following the trend from the best-performing model in the previous experiment, ranging from 168.79 to 2111.71. On the class level, the best-performing models were the ML Objects, ML Object-Captions, and the ML Objects-Human models. However, like the classification model, it is revealed that the models responsible for the best overall and class-level performance are ones that contain ML Objects. Moreover, it is worth highlighting that Asia performs best with the ML Objects model, suggesting that the detected objects in the images from Asia are indicative of income.
The higher performance exhibited by the ML Objects and ML Captions Model is particularly unexpected, especially when contrasted with the trend found in the classification model. However, this observation aligns with the claim made by Zhang et al. (2021) on machines being better at selecting distinguishing features in the context of image segmentation. In our case of income regression, this also applies, although only for ML Objects rather than ML Captions. This could be attributed to certain objects in the images being associated with certain income levels, therefore improving the model’s ability to predict income. Further experiments would need to be conducted to explore this.
Furthermore, we compared the performance differences between the best and worst-performing models (overall), namely the ML Objects-ML Captions Model and the Human Model. The results of the Kolmogorov-Smirnov test showed that the difference is statistically significant ($p$ $<$ 0.001), suggesting that the ML Objects-ML Captions Model’s better performance is statistically reliable.
Figure 4 illustrates a scatter plot of the regression model trained using the ML Objects and ML Captions to predict income on the test dataset. The data points are colour-coded by continent. The blue dotted line represents the line of best fit and the red solid line represents the identity line.
The plot reveals that most of the samples whose incomes are underestimated are mostly countries from Asia, with some samples from the Americas and Europe. On the other hand, most of the overestimated samples are from the Americas, with several samples from Asian and Europe. It is also important to note that none of the African samples deviate significantly from the identity line.
The underestimation of samples from Asia may be attributed to the high number of samples within the lower to middle income range in the training set, resulting in the model skewing predictions for samples from Asia. Interestingly, when looking at the distribution of samples from the Americas, the income levels range from lower to middle income, as well, despite the overestimation by the model. This suggests that perhaps the features detected in the ML Objects and ML Captions for the Americas are not well-represented, causing inaccurate predictions. This is made worse due to the limited number of samples in the dataset, causing there to be less variety in income distribution for particular regions.
Figure 4: Income Regression Model using ML Objects and ML Captions as Annotations Grouped by Continent
Limitations. First, the experiments were done on only one dataset due to the limited availability of image datasets that contain features relating to socio-economic background, such as geographical region and income. Performing the experiments on another similar image dataset would make the results more reliable. Second, the number of image samples in the dataset that have been annotated by humans is rather small. Due to time constraints and limited resources, doing further data collection to obtain more human annotations was not feasible given the scope of this investigation. Having more annotated images would result in better generalisability of the models. Third, due to the small size of the annotated data, under-sampling was not performed for the regression task - unlike the classification model which had under-sampling embedded in the algorithm - which meant that there may be data imbalance when considering number of samples in each region. This could potentially be the reason as to why the performance of some regions is notably better than others. Lastly, there is a risk of misclassification and fault intolerance in this work, as the small dataset could affect the generalisability of the model and, therefore, the trends observed.
Conclusion
Answering the RQs. We first look at RQ1: “How similar are human-generated and ML-generated annotations?”. It can be concluded that from a low-level perspective, i.e., what types of word occur, sentence structures, the ML Objects and ML Captions are similar to each other, but are relatively different to the human labels. This is expected due to underlying cognitive processes during the annotation stage. However, when looking at them from a high-level point of view, despite the apparent low similarity, they are alike in that they exhibit consistent similarity or dissimilarity regardless of the geographical region of the images, indicating lack of regional bias. This led to RQ2: “How do different combinations of annotations affect fairness in ML predictive models?” In the case of region classification, it was revealed that the human labels resulted in the best overall-performing model, as well as the most balanced performance across individual classes. It is additionally apparent that each region obtained best classification performance with models that utilised annotation combinations that included human labels. Similarly, for income regression, the model that has the best overall performance also has the most balanced performance across regions. However, in this case, the combination of ML Objects and ML Captions is what results in this enhanced performance. A similar trend continues in that the best performance for individual classes was obtained with models that contained ML Objects in its annotations. Considering these findings, it can be deduced that there is no particular combination of annotations that result in best overall and most balanced performance as that is highly dependent on the quality of the images and how well the human annotator or machine is able to select the best discriminative features. Additionally, due to imbalanced distributions of income across data samples and low number of samples for certain regions, the regression model is prone to underestimation or overestimation of income for certain regions. Moreover, our analysis suggests that it may be plausible that models with the best overall performance tend to also be more balanced on the class-level, and therefore, fairer. However, further studies need to be done to confirm this.
Implications. These findings have several implications. First, despite their apparent differences, humans and machines may be more alike in their bias, or lack thereof, in the context of visual content. Their consistent similarity and dissimilarity of annotations across regions implies that their level of fairness is comparable. This insight could steer researchers towards further advancing ML models to avoid as much bias as possible. Second, achieving a more balanced performance on the region level can help reduce biases in ML models. This would result in models that not only give accurate outcomes, but ones that are also fair. This would be highly beneficial in real-world scenarios where the model’s outcomes could disproportionately affect certain demographic groups. Third, due to machine annotations (ML Objects and ML Captions) performing better and fairer at the regression task, tasks in other domains could potentially benefit from using machine annotations as well. However, this could result in a potential negative societal impact in that humans become overly reliant on machines, which may cause them to miss possible biases in the models that originated from the training data, miss errors due to assuming the superior performance of the ML models, or even cause difficulties in identifying accountability if the ML models’ prediction errors result in significant repercussions. This could also lead to potential misuse of the technology as it is possible that one intentionally uses a model that is potentially biased without attempting to address the fairness problems.
Acknowledgements
ChatGPT and Claude AI were utilised to assist with brainstorming and to generate sections of this Work, including alternative phrasings and parts of code used for the experiments.

### v2

@@
 We aim to understand differences in perception and identify potential biases in content interpretation.
-Our dataset comprises images of people from various geographical regions and income levels washing their hands.
+Our dataset comprises images of people from various geographical regions and income levels, covering various daily activities and home environments.
 We compare human and ML-generated annotations semantically and evaluate their impact on predictive models.
-Our results show low similarity between human and machine annotations from a low-level perspective, i.e., types of words that appear and sentence structures, but are alike in how similar or dissimilar they perceive images across different regions.
-Additionally, human annotations resulted in best overall and most balanced region classification performance on the class level, while ML Objects and ML Captions performed best for income regression.
-Humans and machines’ similarity in their lack of bias when perceiving images highlights how they are more alike than what was initially perceived.
-The superior and fairer performance of using human annotations for region classification and machine annotations for income regression show how important the quality of the images and the discriminative features in the annotations are.
+Our results show highest similarity between ML captions and human labels from a low-level perspective, i.e., types of words that appear and sentence structures, but all three annotations are alike in how similar or dissimilar they perceive images across different regions.
+Additionally, ML Captions resulted in best overall region classification performance, while ML Objects and ML Captions performed best overall for income regression.
+The varying performance of annotation sets highlights the notion that all annotations are important, and that human-generated annotations are yet to be replaceable.
 Introduction
-Machine Learning (ML) is prevalent in numerous applications, e.g., facial recognition and recommendation systems.
-However, many of these use cases could potentially result in discrimination and unfair decisions.
-For example, it is possible that a facial recognition algorithm is less likely to perform well when predicting people of certain ethnic groups, e.g., people of African heritage may not be identified as accurately as people of European origin.
-This is problematic, especially as we aim to rely more on ML for daily tasks.
-Since supervised ML models require data for training, biased training data could affect the fairness and bias of the learned model – also known as algorithmic bias – which can cause it to generate prejudiced predictions.
+Data annotations are needed to train Machine Learning (ML) models.
+These annotations are usually obtained by having human experts manually label data.
+However, ML models can also be used to produce data annotations, ensuring a more efficient process.
+However, it is unclear what the differences of such annotations would be as compared to human-generated ones.
+Using ML models for data annotation could potentially be problematic if the annotations prove to be unreliable or biased.
+Biased training data could affect the performance and reliability of the learned model – also known as algorithmic bias – which can cause it to generate prejudiced predictions.
 Unfortunately, it is quite likely that the data will contain bias considering that they are commonly annotated by human annotators (Demartini, Roitero, and Mizzaro 2024).
@@
 In this study, we first investigate how similar annotations generated by humans and machines are, and explore possible justifications for their similarities or differences, in the context of how they think and perceive images.
-This leads to our first research question (RQ1): “How similar are annotations generated by humans compared to the annotations generated by ML models?” Moreover, due to the aforementioned biases that could appear in the annotations, we look into how different annotations affect fairness in predictive models.
-This leads to our second research question (RQ2): “How do different combinations of annotations affect fairness in ML predictive models?” To do this, we use the annotations in vector representation to train predictive ML models.
-These annotations are based on images of people washing their hands from varying geographical regions and socio-economic classes.
+This leads to our first research question (RQ1): “How similar are annotations generated by humans compared to the annotations generated by ML models?” Moreover, due to the aforementioned biases that could appear in the annotations, we look into how different annotations affect unbiasedness in predictive models.
+This leads to our second research question (RQ2): “How do different combinations of annotations affect the performance and bias in ML predictive models?” To do this, we use the annotations in vector representation to train predictive ML models.
+These annotations are based on images of people performing different activities from varying geographical regions and socio-economic classes.
 This study contributes to the fields of information retrieval and algorithmic bias in several ways.
-Firstly, it provides a comparative analysis of human-generated and machine-generated annotations in the context of socio-economic diverse content, which give insights on how humans and machines are alike in their bias when perceiving visual content despite their low-level differences, filling a gap in the existing literature.
-Secondly, it explores whether different annotations sets can impact the efficacy and fairness of ensemble-based ML models, thus advancing both theoretical and practical understanding of this critical issue.
+Firstly, it presents a new large-scale crowdsourced annotation dataset for data that captures diverse socio-economic visual content.
+Secondly, it provides a comparative analysis of human-generated and machine-generated annotations in the context of socio-economic diverse content, which gives insights on how humans and machines compare in their bias when perceiving visual content, filling a gap in the existing literature.
+Thirdly, it explores whether different annotations sets can impact the efficacy and bias of predictive ML models, thus advancing both theoretical and practical understanding of this critical issue.
 Related Work
@@
 Devillers, Vidrascu, and Lamel (2005) emphasise how these biases can be dangerous when the training dataset involves actual humans as these learned biases tend to mimic forms of discrimination common amongst humans, such as racial and gender bias.
-An example of this is Microsoft’s “Tay”, an AI chatbot that initially mimicked a teenage girl but became anti-semitic, racist, and sexist over a span of only 16 hours (Zemčík 2020).
 Sun, Nasraoui, and Shafto (2020) further discuss how the problem worsens when users make decisions based on the biased content provided by these models.
@@
 As compared to this body of literature, we focus on comparing human and machine-generated annotations of visual content with the goal of understanding how bias may differ between these annotations.
-Dealing with Imbalanced Data.
-It is important for data to be well-balanced, ensuring that every class in the dataset is well-represented.
-This is especially true for traditional ML models as they were originally designed for datasets that are balanced.
-There are various ML methods that were designed to deal with imbalanced datasets, such as random under-sampling and over-sampling techniques (Estabrooks, Jo, and Japkowicz 2004; Chawla et al.
-2002), creating synthetic data (Chen et al.
-2023; He et al.
-2008; Han, Wang, and Mao 2005), and cost-sensitive learning where the model is penalised more when a minority class sample is incorrectly classified compared to a majority class sample (Ting 2002).
-In our work, we adopt under-sampling techniques to overcome the issue of data imbalance across some classes in our dataset due to the limited number of available samples.
-Ensemble Models.
-Minimising the trade-off between accuracy and fairness is a question that is very relevant to our work.
-There has been research indicating that a loss in accuracy is inevitable when a model is trained with limitations relating to fairness.
-Based on Kenfack et al.
-(2021), we know that compared to individual classifiers, ensemble models have an improved predictive performance, but compromise on fairness.
-In our work, we include the use of ensemble models in our experiment to test whether there is the same effect on our dataset.
+Image Captioning.
+Image captioning has emerged as a critical intersection of computer vision and natural language processing, evolving from template-based approaches to sophisticated deep learning models.
+As Sharma and Padha (2023) outline, this field has found applications ranging from visual aids to medical imaging and security systems, while highlighting ongoing challenges in real-time performance and domain-specific datasets.
+Ghandi, Pourreza, and Mahyar (2023) identify key technical challenges including object hallucination, missing context, and illumination conditions, while Xu et al.
+(2023) frame the captioning process through the stages of seeing, focussing, and telling inspired by humans.
+Kasai et al.
+(2021)’s THumB protocol shows that the quality of captions created by humans substantially surpass those generated by machines in terms of covering key information.
+Moreover, Sarhan and Hegelich (2023) expose critical concerns about AI-generated captions’ tendency toward generic descriptions, which can either erase or stereotype social group identities, particularly in politically sensitive contexts.
+Building on this literature, our work employs the state-of-the-art BLIP model (Li et al.
+2022) with the Vision Transformer architecture (Dosovitskiy et al.
+2020), which has shown to be effective among Vision-Language Pre-training methods, to generate captions that we systematically compare with human annotations across socioeconomic contexts.
 Cognitive Differences in Humans and Machines.
 To this day, human experts are still needed to provide accurate and reliable annotations that could be used to train ML models (Valizadegan, Nguyen, and Hauskrecht 2013).
-Understanding how differently humans and machines think is crucial in helping us develop better and fairer automated systems that could assist us in solving problems, i.e., annotating images for ML.
+Understanding how differently humans and machines think is crucial in helping us develop better and less biased automated systems that could assist us in solving problems, i.e., annotating images for ML.
 Sharifi Noorian et al.
@@
 Methodology
-This section details the methodology we used to address the research questions introduced earlier.
-The methodology begins with data collection from Dollar Street.
-Annotations are preprocessed and added to the data either through crowdsourcing or ML techniques.
-The methodology is divided into two key experiments: the first evaluates annotation similarity; the second investigates the impact of different annotation sets on supervised models, using ensemble learning and comparing the performance of human and machine annotations.
+Figure 1: Handwashing images annotated using the pre-trained Faster R-CNN model.
 Experimental Set-Up
-In this study, we utilised CPU resources to conduct all experiments. The computations were performed locally on a device with an Apple M3 Pro Chip, 11 CPU cores, and 36GB of RAM.
+In this study, we utilised cloud-based computing resources provided by Amazon Web Services.
+The experiments used a mix of instance types, with some CPU-only instances, such as the ml.c6i.16xlarge, and some GPU-enabled instances, such as the ml.r5.16xlarge, and ml.r5.24xlarge instances, to accelerate processing times.
 Dataset
@@
 It features photographs of everyday objects from more than 240 families in 50 countries, providing a representation of living conditions across different economic classes.
-There are 197 unique instances of images and videos utilised in our experiments, consisting of people washing their hands, sourced from various geographical regions and income levels.
-The topic of hand washing was selected due to its crucial role in public health and its practice across all cultures and income levels.
+The dataset contains 3,041 images, sourced from various geographical regions and income levels, consisting of the following image categories: Drinking Water, Drying Clothes, Front Doors, Hand Washing, Kitchens, Living Rooms, Places for Dinner, and Washing Clothes.
+These images include a diverse range of topics, covering both actions and static elements, such as home environments.
+This is to provide a good representation of different aspects of daily life, which is applicable across all cultures and income levels.
 Data Annotation
-Figure 1: Handwashing images annotated using the pre-trained Faster R-CNN model.
-Crowdsourcing.
-The first annotation set is sourced from a structured crowdsourcing process conducted on Amazon Mechanical Turk (MTurk) by Fan et al.
-(2022).
-In their study, crowd workers viewed short videos of individuals washing their hands from various global regions and income levels, providing textual annotations.
-These annotations captured various aspects of handwashing behaviour and the surrounding environmental context.
-The dataset includes annotations for 112 videos, each annotated by 10 US-based workers.
-Fan et al.
-(2022) implemented quality control measures to ensure reliable annotations.
-Consent from the participants was obtained through a signed consent form.
-Additionally, no personally identifiable information or offensive content were collected in the crowdsourcing process.
-We use these existing human annotations in our current study, benefiting from their comprehensive approach to data collection.
 ML Object Detection.
@@
 2017).
-We chose to perform object detection on the dataset as most of the images contain objects, i.e., bottles, sinks, instead of subjects performing actions, i.e., person washing hands.
+We chose to perform object detection on the dataset as most of the images contain objects, e.g., bottles, sinks, instead of subjects performing actions, e.g., person washing hands, despite the fact that some of the image categories are meant to be action-based.
 Due to this, the COCO dataset was chosen as the most suitable for training as it contains annotations of everyday objects.
@@
 We chose to perform image captioning as it would likely resemble the annotations generated by humans.
+Figure 2: Screenshot of the MTurk annotation interface showing the task instructions and image assessment layout.
+Crowdsourcing.
+Our human annotations were collected through crowdsourcing tasks on Amazon Mechanical Turk (MTurk).
+The dataset comprises annotations for 1,886 Dollar Street images covering eight aspects of daily living: drinking water, washing clothes, front doors, hand washing, kitchens, living rooms, places for dinner and washing clothes.
+We recruited MTurk workers to provide annotations for the images.
+To ensure high-quality annotations, we restricted participation to workers with at least 5,000 completed tasks and 98% approval rate.
+Task instructions and interface screenshots are shown in Figure 2.
+Based on our pilot study that measured average task completion times, participants earned an estimated hourly wage of USD 7.5.
+Each image was annotated by three different workers to ensure annotation reliability.
+The total participant compensation for annotating the complete dataset amounted to USD \$373.
+This research was conducted under the approval of Human Research Ethics Committee.
+Participants were informed of minimal risks associated with viewing everyday living condition images and provided informed consent before participation.
+To protect participant privacy and ensure data security, any potentially identifying information was removed from the dataset.
+We also applied post-processing filters to remove low-quality submissions such as random numbers and nonsensical text.
+This approach allowed us to build a comprehensive dataset of human observations across Dollar Street images, which represent a wide range of living conditions worldwide.
 Data Preparation and Preprocessing
-The data cleaning process constituted the initial step in preparing our dataset for analysis.
-In this phase, corrupted images were identified and removed based on criteria including integrity checks and visual inspections.
-Images that did not contain labels from any of the annotation categories are also removed, i.e., images that did not have any objects detected in it.
-This process reduced the original dataset to 197 unique data samples with distinct IDs.
-The annotations were preprocessed to deal with inconsistencies in the human annotations.
+Prior to obtaining all three annotation sets, the images were selected based on whether they successfully led to the generation of ML Objects with at least 50% confidence and ML Captions beforehand.
+Moreover, duplicates were removed and down-sampling was performed to ensure the images were balanced for all regions and income groups.
+The human annotations were processed to deal with inconsistencies.
 These steps involve noise/punctuation removal, conversion to lower case, auto-correcting misspellings, and segmenting strings that are concatenated.
-We utilised a sentence embedding model to convert the three types of annotations into sentence embeddings.
-Instead of using a pre-trained sentence transformer model to generate the embeddings, we trained a sentence transformer model from scratch on all three sets of annotations.
-We did this as we want the model and representations to be very focused, and due to the limited number of samples, fine-tuning an existing sentence transformer, such as sBERT, resulted in worse performance.
-To create the sentence transformer, we used a pre-trained Transformer model, “distilroberta-base”, which is available at the Hugging Face Hub (Sanh et al.
-2019), to obtain the word embedding model.
-We then combined this word embedding model and a mean pooling function over the token embeddings so that it could generate a fixed-size vector as output.
-Using the sentence transformer model, we generated embeddings with a size of 384 dimensions.
-Experiment 1: Investigating Annotation Similarity (RQ1)
+The number of images in the dataset decreased at each step, starting at 3,041 images at the start, reducing to 2,797 after generating the ML Objects and Captions, then to 1,888 after removing duplicates and down-sampling, and finally to 1,886 unique samples after crowdsourcing.
+In total, 1,155 samples were discarded from the original dataset.
+We utilised a fine-tuned sentence transformer model to convert the three types of annotations into sentence embeddings.
+We opted to fine-tune an existing model instead of using an pre-trained model out of the box as we want the model and representations to be very focused to our data.
+We fine-tuned a pre-trained Sentence Transformer model, “all-MiniLM-L12-v2”, available at the Hugging Face Hub [^1] https://huggingface.co/sentence-transformers/all-MiniLM-L12-v2.
+The fine-tuning process was run over 5 epochs with a batch size of 16.
+The model was fine-tuned on the three annotation sets so that it can better capture the characteristics of the annotations, hence generating closer embeddings.
+The generated embeddings have a size of 384 dimensions, which is the default embedding size from the original pre-trained model.
+Exp. 1: Investigating Annotation Similarity (RQ1)
 To investigate the similarity between human, machine-generated object labels, and machine-generated captions, we utilised the embeddings that have been generated for the three types of annotations across all images.
@@
 Doing this helps us analyse the distances of the vectors, and thus the annotations, for every image in the dataset and see how they are distributed.
-We then used the sentence transformer model’s built-in similarity function to compute the cosine similarity scores between each possible pair of annotations.
+We then used the sentence transformer model’s built-in cosine similarity function to compute the similarity scores between each possible pair of vector embeddings representing the annotations.
 Using these scores, we look at the average similarity scores for each pair and see how the pairwise scores are distributed using scatterplots and boxplots.
-These experiments are appropriate for addressing RQ1 as t-SNE and computation of similarity scores across pairs allows us to evaluate how similar the annotations are, whereas the pairwise scatter plot helps us derive trends and patterns between annotations, further evaluating their likeness to one another.
-Experiment 2: Exploring How Annotations Impact Fairness in Predictive Models (RQ2)
+We computed descriptive statistics about the length of the annotations to help us better understand differences between the sets.
+We compared the average number of words each worker inputted during crowdsourcing to the average number of words in each caption, as well as the average number of object labels predicted in each image.
+These experiments are appropriate for addressing RQ1 as t-SNE, computation of similarity scores across pairs, and annotation length allow us to evaluate how similar the annotations are, whereas the pairwise scatter plot helps us derive trends and patterns between annotations, further evaluating their likeness to one another.
+Exp. 2: Exploring How Annotations Impact Performance and Bias in Predictive Models (RQ2)
+This section emphasises the diversity of model prediction outcomes over overall model effectiveness.
+These experiments aim to evaluate the distribution of prediction errors across different regions and categories, allowing us to identify potential discrimination, e.g., higher prediction errors for certain subsets of the population as compared to the error on others.
 To investigate how the three annotation sets affect predictive models, we use the vector embeddings of every annotation as training data for two supervised learning models.
+As previously mentioned, the dataset was balanced for all geographical regions and income groups, as they are the target variables for these predictive tasks.
+Since ‘image category’ is not a feature in the training data (annotation embeddings) and is only used in a group-by operation after model training, the metric scores for each category are independent of whether ‘image category’ was balanced.
 Before training the models, the dataset was split into training and test sets with a split ratio of 80:20, respectively.
@@
 Grid Search CV was performed in every independent run.
-The first model is an RUS Boost Classifier used to classify each sample on region (Seiffert et al.
+The first model is an RUS Boost Classifier used to classify annotation embeddings on geographical regions, which originate from the dataset, namely Africa, Asia, Europe, and the Americas (Seiffert et al.
 2008).
-This classifier involves the incorporation of the Adaptive Boost algorithm, which is a sequential ensemble technique that can lower a model’s bias and variance, with random under-sampling, which helps deal with data imbalance (Mohammed and Kora 2023; Seiffert et al.
+This classifier involves the incorporation of the Adaptive Boost algorithm, which is a sequential ensemble technique that can lower a model’s bias and variance (Mohammed and Kora 2023; Seiffert et al.
 2008).
-This makes the algorithm ideal for our dataset due to its small size and there not being an equal number of samples from every region.
 The base estimator used in the model is the Decision Tree Classifier (Swain and Hauska 1977).
@@
 The trained classifier is run 10 times with different random seeds to obtain a mean F1 score for each class (region) and a mean overall weighted F1 score, with their respective standard deviations.
-This is done to evaluate the model effectiveness for each class and the overall model performance.
+This is done to evaluate the overall model performance, as well as its performance when grouped by geographical region and image category.
 The model is run multiple times to allow us to evaluate its stability.
-The Welch’s t-test was performed to compare whether the difference in performance of the best and worst overall-performing models are statistically significant.
+The Welch’s t-test was performed to compare whether the difference in performance of the best and worst overall-performing models is statistically significant.
 This test was performed due to the Gaussian distribution of the F1-Scores and varying variances between the models.
-The second model uses the Adaboost.R2 algorithm to predict each sample’s income (Drucker 1997).
+The second model uses the Adaboost.R2 algorithm to predict income based on the annotation embeddings (Drucker 1997).
+Income was chosen as a target variable as it is one of the labels associated with each image in the dataset.
 Similar to the region classifier, this algorithm uses a Decision Tree Regressor as its base estimator.
-Since boosting helps reduce prediction error on the small dataset, this algorithm is suitable for this task.
+Since boosting helps reduce prediction error, this algorithm is ideal for this task.
 Using grid search CV, we optimised for the number of estimators, learning rate, and the type of loss function used in the Adaboost.R2 algorithm.
 The model was executed 10 times with different random seeds.
-The root mean square error (RMSE) was obtained for every run, both for overall and region-based evaluation.
-The RMSE takes the square root of the mean square error, which performs the measurement using the same units as the target variable, i.e., income level.
-A regression plot is generated to visualise the distribution of predicted samples by income for the best-performing model.
+The root mean square error (RMSE) was obtained for every run to allow for evaluation of the model performance, including overall, by region, and by category.
+A regression plot was generated to visualise the distribution of predicted samples by income for the best-performing model.
 The Kolmogorov-Smirnov test was performed to compare whether the difference in performance of the best and worst overall-performing regression models are statistically significant.
 This test was performed due to the non-Gaussian distribution of the RMSE scores between the models.
-These predictive models are appropriate in addressing RQ2 as it helps us investigate how often each region is correctly identified for a given image, and how well income is predicted for images in a given region.
-For evaluation purposes, the F1-score is appropriate and sufficient as it is a good metric for evaluating classification performance on imbalanced datasets, as ours is, whereas the RMSE scores are appropriate and sufficient as the measurement being done in the same units as the target variable makes it easier to interpret.
-Looking at these evaluation metrics for every region helps us determine whether the models perform better for particular regions, giving insight on how fair the model is.
+These predictive models are appropriate in addressing RQ2 as they help us investigate how well region and income are correctly identified for a given image annotation set.
+For evaluation, the F1-score is appropriate and sufficient ideal as it is a robust metric which balances precision and recall.
+The RMSE scores are appropriate and sufficient as the measurement being done in the same units as the target variable makes it easier to interpret.
+Looking at these evaluation metrics for every region and category helps us determine whether the models perform better for particular regions and categories, giving insight on how biased the model is.
 Results and Discussion
@@
 Figure 3: t-SNE visualisation of sentence embeddings of each label
-Annotations Used Mean ± s.d.
-ML Objects and ML Captions 0.87 ± 0.05
-ML Objects and Human Labels 0.42 ± 0.1
-ML Captions and Human Labels 0.45 ± 0.06
-Table 1: Mean and standard deviation of similarity scores between each pair of annotations across all samples.
-Figure 3: A plot of pairwise relationships between each pair’s similarity score.
 Figure 3 illustrates a t-SNE visualisation of the vector embeddings of the three annotation categories for each data sample.
-The results show how, despite having trained an embedding model from scratch, the embeddings still form into three distinct clusters.
-Further analysis shows that the human labels cluster is furthest from the ML captions cluster, with the ML object labels cluster in between them.
+The results show how, despite having a fine-tuned embedding model, the embeddings form into two distinct clusters, with one for ML Captions and Human labels and another for ML Objects.
+This evidence implies that there exist differences between the different types of annotations we collected in the study.
 The clusters of embeddings are due to differences in the annotations.
-Human annotations are holistic, encapsulating both the objects and the array of actions unfolding in the captured images and videos.
+Human annotations are holistic, encapsulating both the objects and the array of actions unfolding in the captured images.
 On the other hand, the ML-generated object labels are primarily object-centric.
 The COCO dataset, which the ML object detection model is trained on, does not extend to action recognition, which is not relevant due to the object-oriented images in the dataset, thereby inherently limiting the ML-generated object annotations to object identification.
-The ML-generated captions, although more descriptive than the object labels, tend to not be as specific as the human annotations since they are more likely to contain easily recognisable actions and objects rather than complex descriptions of people’s emotions or the environment.
-Additionally, there are apparent syntactic differences between the annotations - the ML object labels consists of lists of objects per image; the ML captions consist of a phrase per image; while the human labels are a list of words and phrases.
-It is also important to note that a limitation that could have affected this distinct categorisation is that stop word removal and lemmatisation were not performed during preprocessing as to preserve nuances and context that contain relevant information.
+The ML-generated captions would be more descriptive than the object labels and although they may not be as specific as the human annotations, they are more likely to contain actions and objects, which may also be present in Human Labels.
+Additionally, there are apparent syntactic differences between the annotations - the ML Objects consists of lists of objects per image; the ML Captions consist of a phrase per image; while the Human Labels are a list of words and phrases.
+It is also important to note that a limitation that could have affected this categorisation is that stop word removal and lemmatisation were not performed during preprocessing as to preserve nuances and context that contain relevant information.
 This causes the sentence structures in the annotations to remain unchanged, resulting in more pronounced categorisations.
-Considering these differences, the distinct clusters in the visualisation are expected.
-However, since t-SNE is only a two-dimensional representation of the data, it would be useful to also look at the similarity scores, generated by the BLIP model’s cosine similarity function, between each pair of annotations across the dataset.
+Considering these differences, the clusters in the visualisation are expected.
+Annotations Used Mean ± s.d.
+ML Objects and ML Captions 0.50 ± 0.12
+ML Objects and Human Labels 0.49 ± 0.12
+ML Captions and Human Labels 0.69 ± 0.12
+Table 1: Mean and standard deviation of similarity scores between each pair of annotations across all samples.
+Since t-SNE is only a two-dimensional representation of the data, it is useful to also look at the similarity scores, generated by the BLIP model’s cosine similarity function, between each pair of annotations across the dataset.
 This is shown in Table 1.
 These values give a more accurate and quantitative comparison of the annotation similarities.
-It can be seen that the ML object labels and the ML captions are the most similar pair of annotations (0.87), followed by ML captions and human labels (0.45), and ML object labels and human labels (0.42).
-It is important to note that when comparing human labels with ML-generated ones, the similarity scores are relatively low, regardless of the type of ML annotation.
-The high similarity of both ML annotation types is a reasonable outcome.
-Although the annotations were generated by two different pre-trained ML models which were trained with different types of training data, i.e., labels versus image-text pairs, the training data both originated from the COCO dataset and are based on the same set of images.
-This means that there would be more overlap of objects between the two annotation categories.
-In contrast, the low similarity values between either ML annotations with the human labels is expected since there may be higher variety of words and phrases that were used by the human annotators.
-Moreover, when we compare the t-SNE visualisation and the similarity scores, it may seem inconsistent at first glance due to the ML objects and human labels pair having the lowest similarity score despite being seemingly furthest apart visually.
-However, since t-SNE is a dimensionality reduction technique, there may be features that were lost in the process, resulting in less accurate representations.
+It can be seen that the ML Captions and Human Labels are the most similar pair of annotations (0.69), followed by ML Objects and ML Captions (0.50), and ML Objects and Human Labels (0.42).
+It is important to note that when comparing ML Objects to any of the other annotations, the similarity scores are relatively low.
+To further investigate length statistics per image, Human Labels averaged 7.18 words per worker, ML Captions averaged 8.62 words per caption, and ML Objects averaged 4.36 object labels per prediction.
+The length of ML Objects is determined by the number of objects in the predictions rather than the average number of words as ML Objects consists of a discrete list of items.
+Figure 4: Pairwise relationships between similarity scores.
+The higher similarity between ML Captions and Human Labels can be explained by several factors.
+First, both humans and ML captioning models can recognise not only objects but also actions in images, leading to overlapping features in their descriptions.
+Additionally, the likelihood of descriptive phrases appearing in both categories is much higher, which contributes to their high similarity.
+This is also further supported by the t-SNE visualisation which shows both of these annotations placed relatively close to each other in the same cluster.
+In contrast, ML Objects having a low similarity to both ML Captions and Human Labels is expected as ML Objects annotations are more likely to be syntactically different and may have a more limited range of vocabulary compared to the other annotation types.
+Moreover, the average length of the annotations shows that ML Captions and Human Labels have similar average lengths (8.62 and 7.18, respectively) compared to ML Objects (4.36).
+This further supports the interpretation that ML Captions are similar to Human Labels due to their sentence structures.
 To further explore the relationships of these similarity scores, a plot of pairwise relationships is presented in figure 4.
@@
 The plot is colour-coded by region to allow further analysis of similarity scores by geographical location.
-The analysis reveals that there are no particular trends or patterns with respect to region.
-Moreover, when looking at the subplots for ML objects-ML captions pairs, i.e., subplots on first column/row, the similarity scores tend to be relatively high, on average, with the exception of several outliers.
-This aligns with the previous observation that ML objects and ML captions have the highest average similarity score (0.87).
-This pair also has the smallest variance out of all pairs, which is also evident when comparing the vertical spread of scores across the subplots in all rows, i.e., the scores of the ML objects-ML captions pair are not spread as much vertically when disregarding the outliers.
-In contrast, the ML objects-human labels pair, i.e., second row/column, appears to have the highest variance, which also aligns with the standard deviation of the mean similarity score observed in the previous experiment (0.1).
-It is also shown that the ML objects-ML captions pair has the same outliers as the ML objects-human labels pair (first row, second column), but not with the ML captions-human labels pair (first row, third column).
-Interestingly, the ML captions-human labels pair (third column/row) seems to have no outliers.
-The lack of discernible patterns in the way the scores are distributed across regions suggests that the images have little to no effect on how similar two sets of annotations will be.
-This could mean that the discriminative features in the images, or lack thereof, are detected equally by both humans and ML models.
-Therefore, none of the annotation sets describe images from a particular region any better or worse than the other.
-This also suggests that humans and machines may be similar in terms of how biased they are when perceiving images.
-The low variance of similarity scores for the ML objects-ML captions pair suggests that for the majority of images, both annotations consistently contain similar objects, which - similar to the previous discussion on their high similarity scores - is expected due to the origin of the both models’ training data.
-Similarly, the high variance of scores for the ML objects-human labels pair aligns with expectations.
-This is due to, not only syntactic differences, i.e., objects vs longer phrases or sentences, but also differences in the way machines and humans interpret visual content, i.e., machines are more likely to capture easily recognisable objects and actions but may fail to capture more complex and nuanced observations, such as environment aesthetics or partially obstructed objects.
-Moreover, the fact that the ML objects-ML captions pair has the same outliers as the ML objects-human labels pair only (first row, second column) indicates that for those particular images, the ML object detection model may be detecting objects that do not appear in the corresponding ML captions or human labels, or failing to identify objects that do.
+The analysis reveals that most regions show no particular trends or patterns, with one exception: images from Europe tend to cluster in the higher similarity range (top right area) of the plots, indicating stronger agreement between all annotation types for European images.
+Moreover, when looking at the subplots for ML Captions-Human Labels pairs, i.e., subplots on the third column/row, the similarity scores tend to be relatively high, on average, with the exception of several outliers.
+This aligns with the previous observation that ML Captions and Human Labels have the highest average similarity score (0.69).
+It can also be seen that all three pairs show the same amount of spread of scores across all subplots, which correspond to the variance values of the mean similarity scores between each pair (0.12).
+The lack of discernible patterns in the way the scores are distributed across the Americas, Africa, and Asia suggests that images from these regions have little to no effect on how similar two sets of annotations will be.
+This could mean that the discriminative features in these regions’ images, or lack thereof, are detected equally by both humans and ML models.
+Therefore, none of the annotation sets describe images from these three regions any better or worse than the other.
+On the other hand, the cluster of European samples towards the top right of the plots suggests that European annotations from each group seem to share strong similarities with every other annotation type.
+This may mean that both humans and machines were able to perform more consistently across many European images compared to non-European ones.
+We hypothesise that this due to the fact that both humans and machine-based models are more familiar with Western societies images based on their past experience or training.
+Moreover, the fact that the ML Objects annotations have a relatively low similarity score indicates that the ML object detection model may be detecting objects that do not appear in the corresponding ML Captions or Human Labels, or failing to identify objects that do.
 This reasoning is consistent with the findings of Zhang et al.
@@
 The ML captioning model as well as human annotators have the advantage in that they are able to capture non-object elements, such as actions, e.g., “wash”, or environmental descriptions, e.g., “dirty”.
-RQ2: How do different combinations of annotations affect fairness in ML predictive models?
-F1-Score (Mean ± Standard Deviation)
-Annotations Used Africa Asia Europe The Americas Overall
-ML Objects 0.43 ± 0.05 0.45 ± 0.06 0.30 ± 0.09 0.51 ± 0.10 0.42 ± 0.03
-ML Captions 0.60 ± 0.05 0.32 ± 0.05 0.57 ± 0.02 0.38 ± 0.06 0.45 ± 0.02
-Human Labels 0.62 ± 0.05 0.54 ± 0.13 0.57 ± 0.05 0.51 ± 0.11 0.56 ± 0.07
-ML Objects & ML Captions 0.64 ± 0.07 0.50 ± 0.05 0.49 ± 0.06 0.46 ± 0.09 0.53 ± 0.04
-ML Objects & Human Labels 0.57 ± 0.05 0.41 ± 0.05 0.57 ± 0.06 0.55 ± 0.10 0.51 ± 0.04
-ML Captions & Human Labels 0.60 ± 0.09 0.38 ± 0.05 0.62 ± 0.07 0.51 ± 0.12 0.51 ± 0.03
-ML Objects & ML Captions & Human Labels 0.65 ± 0.06 0.46 ± 0.04 0.51 ± 0.06 0.53 ± 0.11 0.53 ± 0.05
-Table 2: Comparison of the region classification model performance across different annotation types.
-The mean and standard deviation of the F1-Score were recorded over 10 independent runs, both for the overall model performance and for each class individually.
-RMSE (Mean ± Standard Deviation)
-Annotations Used Africa Asia Europe The Americas Overall
-ML Obj 256.17 ± 88.2 1161.35 ± 116.23 2467.03 ± 250.75 1142.27 ± 83.42 1341.65 ± 85.56
-ML Captions 179.69 ± 54.86 1456.47 ± 132.90 2164.40 ± 99.97 1815.89 ± 59.91 1564.56 ± 48.54
-Human 228.35 ± 55.14 1657.22 ± 107.99 2675.59 ± 104.75 1433.51 ± 74.91 1633.25 ± 49.57
-ML Obj & ML Captions 168.79 ± 39.57 1223.70 ± 137.91 2111.71 ± 138.83 1191.59 ± 136.96 1279.39 ± 57.66
-ML Objects & Human 117.97 ± 11.04 1497.18 ± 86.39 2621.28 ± 219.83 1108.55 ± 85.18 1480.72 ± 78.63
-ML Captions & Human 149.31 ± 16.17 1523.02 ± 108.62 2525.53 ± 186.66 1202.37 ± 72.59 1486.67 ± 86.93
-Obj & Captions & Human 125.85 ± 8.27 1407.31 ± 54.97 2549.58 ± 135.91 1191.46 ± 82.55 1449.52 ± 45.33
-Table 3: Comparison of the income regression model performance across different annotation types.
-The mean and standard deviation of the RMSE were recorded over 10 independent runs, both for the overall model performance and for each class individually.
+RQ2: How do different combinations of annotations affect the performance and bias in ML predictive models?
+Annotations Used F1-Score (Mean ± Standard Deviation)
+Africa (471) Asia (471) Europe (472) The Americas (472) Overall (1,886)
+ML Objects 0.46 ± 0.02 0.29 ± 0.02 0.41 ± 0.02 0.31 ± 0.03 0.37 ± 0.01
+ML Captions 0.48 ± 0.02 0.35 ± 0.04 0.48 ± 0.02 0.34 ±0.04 0.41 ± 0.01
+Human Labels 0.42 ± 0.02 0.32 ± 0.01 0.38 ± 0.02 0.29 ± 0.02 0.35 ± 0.01
+ML Obj & ML Capt 0.42 ± 0.04 0.32 ± 0.03 0.47 ± 0.01 0.31 ± 0.03 0.38 ± 0.01
+ML Obj & Human Labels 0.43 ± 0.01 0.32 ± 0.02 0.41 ± 0.02 0.29 ± 0.03 0.36 ± 0.01
+ML Capt & Human Labels 0.42 ± 0.02 0.30 ± 0.02 0.40 ± 0.03 0.31 ± 0.02 0.36 ± 0.01
+ML Obj, ML Capt, Human 0.50 ± 0.01 0.31 ± 0.02 0.44 ± 0.02 0.30 ± 0.02 0.39 ± 0.01
+Table 2: Comparison of the region classification model performance across different annotation types grouped by region (target variable), along with the number of samples per region and the overall total.
+Over 10 independent runs, the mean and standard deviation of the F1-Scores were recorded, both for the overall performance and each class (region) individually.
+Annotations Used F1-Score (Mean ± Standard Deviation)
+Drinking
+Water
+Drying
+Clothes
+Front
+Doors
+Hand
+Washing
+Kitchens
+Living
+Rooms
+Places
+for
+Dinner
+Washing
+Clothes
+Overall
+(1,886)
+MLO 0.35 ± 0.03 0.29 ± 0.07 0.28 ± 0.03 0.39 ± 0.03 0.44 ± 0.04 0.30 ± 0.03 0.40 ± 0.04 0.40 ± 0.04 0.37 ± 0.01
+MLC 0.42 ± 0.04 0.38 ± 0.07 0.40 ± 0.05 0.50 ± 0.03 0.39 ± 0.04 0.37 ± 0.03 0.37 ± 0.04 0.40 ± 0.06 0.41 ± 0.01
+H 0.32 ± 0.05 0.29 ± 0.04 0.18 ± 0.02 0.41 ± 0.02 0.45 ± 0.02 0.34 ± 0.03 0.28 ± 0.03 0.43 ± 0.02 0.35 ± 0.01
+MLO & MLC 0.38 ± 0.04 0.45 ± 0.04 0.22 ± 0.03 0.48 ± 0.05 0.48 ± 0.03 0.27 ± 0.03 0.34 ± 0.02 0.39 ± 0.07 0.38 ± 0.01
+MLO & H 0.38 ± 0.03 0.19 ± 0.03 0.30 ± 0.03 0.30 ± 0.02 0.40 ± 0.03 0.38 ± 0.02 0.43 ± 0.03 0.37 ± 0.03 0.36 ± 0.01
+MLC & H 0.38 ± 0.02 0.45 ± 0.05 0.23 ± 0.03 0.38 ± 0.03 0.47 ± 0.03 0.20 ± 0.02 0.30 ± 0.04 0.43 ± 0.03 0.36 ± 0.01
+MLO, MLC, H 0.32 ± 0.03 0.40 ± 0.04 0.29 ± 0.03 0.42 ± 0.02 0.45 ± 0.02 0.33 ± 0.02 0.39 ± 0.03 0.43 ± 0.04 0.39 ± 0.01
+Table 3: Comparison of the region classification model performance across different annotation types grouped by image category (image category was not included in training), along with the number of samples per category and the overall total.
+Over 10 independent runs, the mean and standard deviation of the F1-Scores were recorded, both for the overall model performance and each image category individually.
+Annotations Used RMSE (Mean ± Standard Deviation)
+Africa (471) Asia (471) Europe (472) The Americas (472) Overall (1,886)
+ML Objects 1099.26 ± 106.03 1578.72 ± 91.81 2935.26 ± 85.79 1742.49 ± 55.11 1979.1 ± 45.47
+ML Captions 601.73 ± 25.31 1634.27 ± 12.86 2973.81 ± 57.28 1620.76 ± 26.25 1925.1 ± 26.13
+Human Labels 641.82 ± 30.95 1408.57 ± 35.75 2880.65 ± 24.67 1600.72 ± 29.54 1842.57 ± 14.99
+ML Obj & ML Capt 622.80 ± 63.79 1468.67 ± 26.75 2777.27 ± 83.50 1624.42 ± 54.18 1817.49 ± 37.13
+ML Obj & Human Labels 581.71 ± 28.81 1515.85 ± 164.94 2895.28 ± 31.16 1620.09 ± 35.57 1870.01 ± 41.36
+ML Capt & Human Labels 569.12 ± 12.92 1420.02 ± 31.07 3297.25 ± 145.16 1578.29 ± 35.61 2006.36 ± 58.58
+ML Obj, ML Capt, Human 561.42 ± 29.05 1439.70 ± 40.32 2890.99 ± 53.85 1585.46 ± 41.64 1842.89 ± 30.83
+Table 4: Comparison of the income regression model performance across different annotation types grouped by region (target variable), along with the number of samples per region and the overall total.
+Over 10 independent runs, the mean and standard deviation of the RMSE were recorded, both for the overall model performance and each class (region) individually.
+Annotations Used RMSE
+Drinking
+Water
+Drying
+Clothes
+Front
+Doors
+Hand
+Washing
+Kitchens
+Living
+Rooms
+Places
+for
+Dinner
+Washing
+Clothes
+Overall
+(1,886)
+MLO 1738.35 1699.04 3026.46 1418.68 2502.32 1602.33 1421.24 2030.24 1979.10
+±59.40 ±119.83 ±75.53 ±87.05 ±66.97 ±71.54 ±119.54 ±73.80 ±45.47
+MLC 1765.65 1241.21 2864.27 1197.36 2389.16 1709.97 1593.86 2112.19 1925.10
+±34.95 ±44.87 ±44.27 ±35.92 ±14.50 ±232.10 ±26.17 ±49.54 ±26.13
+H 1571.56 1342.43 2744.72 1363.50 2365.00 1483.01 1242.21 2143.17 1842.57
+±23.31 ±70.87 ±40.38 ±42.24 ±56.15 ±55.17 ±35.88 ±41.70 ±14.99
+MLO & MLC 1531.87 1191.23 2758.96 1376.72 2354.67 1497.12 1341.66 2001.87 1817.49
+±54.58 ±67.30 ±54.13 ±56.35 ±127.14 ±76.62 ±46.02 ±54.91 ±37.13
+MLO & H 1545.01 1399.87 2729.15 1343.64 2385.23 1515.60 1412.33 2175.34 1870.01
+±48.62 ±$4$26.66 ±30.48 ±43.41 ±23.03 ±293.80 ±41.00 ±41.36
+MLC & H 1590.08 1246.10 2771.16 1258.17 2423.63 2818.41 1337.43 2057.95 2006.36
+±49.66 ±29.81 ±20.11 ±19.13 ±55.94 ±407.44 ±32.27 ±26.72 ±58.58
+MLO, MLC, H 1536.57 1250.52 2761.17 1294.69 2373.87 1538.69 1327.09 2143.24 1842.89
+±48.04 ±36.12 ±20.38 ±42.16 ±94.76 ±109.56 ±70.26 ±55.33 ±30.83
+Table 5: Comparison of the income regression model performance across different annotation types grouped by image category (image category was not included in training), along with the number of samples per category and the overall total.
+Over 10 independent runs, the mean and standard deviation of the RMSE were recorded, both for the overall model performance and each image category individually.
 This subsection aims to compare the efficacy of various annotation combinations - ML Objects, ML Captions, and Human Labels - to enhance the performance of models devised for geographical region classification and income regression.
 The performance of the models was appraised using the F1 score, which balances precision and recall, and Root Mean Square Error (RMSE), which measures performance using the same units as the target variable.
-Table 2 tabulates the F1-Scores of the region classification ensemble model across various combinations of the three annotation sets: ML Objects, ML Captions, and Human Labels.
-The mean and standard deviation of the F1-Scores across 10 independent runs were recorded, both for the overall model and for each class (region).
-It is discerned that the Human Labels resulted in the highest overall model performance (0.56).
-The models utilising combinations of the human and ML annotations exhibited poorer overall performance, followed by the ML Captions Model (0.45) and the ML Objects Model (0.42).
-Compared to other models, the Human Labels model has the smallest range of F1-Scores across all classes, ranging from 0.51 to 0.62.
-Moreover, when looking at per-class performance, it is revealed that each class best performs with different models.
-Africa (class) exhibits the highest performance with the model that uses all three annotation sets (0.65), Asia with the Human Labels Model (0.54), Europe with the ML Captions and Human Labels Model (0.62), and the Americas with the ML Objects and Human Labels Model (0.55).
-One notable observation is that regardless of overall or per-class performance, human labels are always used in the models - none of the ML-only models resulted in superior overall or per-class performance.
-The superior overall performance of the Human Labels model is expected.
+Table 2 shows the F1-Scores of the region classification ensemble model across various combinations of the three annotation sets: ML Objects, ML Captions, and Human Labels.
+Table 5 shows the same F1-Scores, but is grouped into the different image categories in the data.
+The mean and standard deviation of the F1-Scores across 10 independent runs were recorded.
+We observe that ML Captions resulted in the highest overall model performance (0.41).
+The model utilising all three annotation sets exhibited slightly poorer overall performance (0.39), followed by the ML Objects and ML Captions Model (0.38), the ML Objects Model (0.37), the models utilising combinations of the Human Labels with either of the other two ML-based annotations (both at 0.36), and the Humans Label Model (0.35).
+Moreover, when looking at per-class performance, we can see that the ML Captions Model performs best for Asia, Europe, and the Americas, at 0.35, 0.48, and 0.34, respectively, whereas the model utilising a combination of all three sets performs best for Africa, showing possible discrimination of ML Captions towards scenes based in the African continent.
+One notable observation is that regardless of overall or per-class performance, ML Captions are always used in the models, whether it would be as a stand-alone or in combination with other annotations.
+When looking through the per-category performance, it is evident that although the ML Captions Model continues to be the best-performing for the Drinking Water, Front Doors, and Hand Washing categories, all seven other categories are better suited with the other annotation models.
+The model containing both ML Objects and ML Captions was the most effective for the Drying Clothes and Kitchens categories.
+ML Objects and Human Labels excelled in handling the Living Rooms and Places for Dinner categories.
+ML Captions and Human Labels worked best for Drying Clothes and Washing Clothes.
+Moreover, Washing Clothes also achieved optimal performance with the Human Labels Model and the all-annotations model.
+It is also important to note that the best-performing models in the Washing Clothes category are ones that contain Human Labels.
+It is also the only category to have the Human Labels Model as its best-performing model.
+Similarly, Drying Clothes, which also has more than one optimal model, achieve best performance with annotation sets that contain ML Captions.
+The superior overall performance of the ML Captions Model is expected.
 As discussed by Zhang et al.
-(2021), humans are more likely to use contextual reasoning in their decision-making process.
+(2021), humans are not necessarily better than machines in choosing discriminative features.
 Although the paper talks about this in the context of humans segmenting images, this may similarly extend to annotation tasks, positively impacting the model’s performance in classifying regions.
-Perhaps it is the case that classifying regions in particular requires more than descriptions of objects and actions, both of which are what the ML models were trained on.
-This could likewise influence the fairer and more balanced class-level performance of the Human Labels model, as evidenced by the range of the F1-Scores across all classes.
-This may also account for why the best-performing models, both overall and class-level, are those that involve human labels.
-While Zhang et al.
-(2021) claim that humans are not necessarily better than machines in choosing discriminative features for image segmentation, our classification model results highlight a different perspective.
-It is evident that the usage of human labels, whether on its own or in combination with another annotation set, leads to more accurate region classifications.
-From this, it can be deduced that humans are, in fact, better than machines at creating image annotations that contain these discriminative features in the context of region classification.
-Using the Welch’s t-test, we found that there are statistically significant differences between the overall F1-scores from 10 independent runs across the Human Labels model and the ML Objects model at $p$-value $<$ 0.001.
-Table 3 shows the Root Mean Square Error (RMSE) for the AdaBoost Regression algorithm used to predict income.
+It is evident that the usage of ML Captions on their own leads to more accurate region classifications for Asia, Europe, and the Americas.
+This suggests that images from these regions may have discriminative features, such as objects and actions, that are captured by the ML Captions and reflected in the model’s predictions.
+Although the ML Captions Model worked quite well for African samples as well (0.48), combining all annotations seemed to perform better for this region.
+This suggests that there may be certain features that the ML Captions Model fails to capture on its own.
+It may be that there are objects that were detected by the ML Objects annotations that were not detected by the ML Captions, as captions tend to convey the overall meaning of the image and overlook objects in the background or unrelated to the main theme of the image.
+Additionally, the fact that ML Captions are always used in the best-performing models shows how the features captured by the captions tend to be very indicative of what region the image is from.
+The variation in optimal models across the image categories is understandable.
+Images from each of the categories may contain varying objects and/or actions that are associated with certain geographical regions, e.g., for theHand Washing category, images of buckets may appear more frequently in samples from Africa, whereas images of sinks may appear more in samples from Europe.
+This explains why certain annotations are better than others at capturing relationships between the objects or actions and the regions.
+The best-performing models for five of the categories (Drying Clothes, Kitchens, Living Rooms, Places for Dinner, and Washing Clothes) are the ones with a combination of at least two annotation sets.
+This may be due to how the images from these categories may have discriminative features that cannot be captured with just one annotation set.
+Using the Welch’s t-test, we found that there are statistically significant differences between the overall F1-scores from 10 independent runs across the ML Captions Model and the Human Labels Model at $p$-value $<$ $1.15\times 10^{-7}$.
+This statistically significant value suggests that the ML Captions Model’s better performance is statistically reliable.
+Table 4 shows the Root Mean Square Error (RMSE) for the AdaBoost Regression algorithm used to predict income grouped by region, while Table 5 shows the scores grouped by image category.
 Similar to the classification model, the experiment was conducted over 10 independent runs across different combinations of the three annotation sets.
@@
 It is apparent that, unlike the classification model, the regression model with ML Objects and ML Captions performed the best overall, indicated by its lowest RMSE value (1817.49).
-This model also has the smallest range of RMSE values across all classes, following the trend from the best-performing model in the previous experiment, ranging from 168.79 to 2111.71.
-On the class level, the best-performing models were the ML Objects, ML Object-Captions, and the ML Objects-Human models.
-However, like the classification model, it is revealed that the models responsible for the best overall and class-level performance are ones that contain ML Objects.
-Moreover, it is worth highlighting that Asia performs best with the ML Objects model, suggesting that the detected objects in the images from Asia are indicative of income.
-The higher performance exhibited by the ML Objects and ML Captions Model is particularly unexpected, especially when contrasted with the trend found in the classification model.
-However, this observation aligns with the claim made by Zhang et al.
+The best-performing models differ across all regions.
+The model containing all annotations was the most effective for African images (561.42), similar to the classification model, while the Asian images achieved the best results with the Human Labels Model (1408).
+Europe saw optimal performance from the ML Objects and ML Captions Model (2777.27), and the Americas was best handled by the ML Captions and Human Labels Model (1578.29).
+Moreover, it is worth highlighting that Asia is the only region where a stand-alone annotation set achieved the best performance.
+Figure 5: Income Regression Model using ML Objects and Captions as Annotations Grouped by Continent.
+The x-axis represents the ground truth values, plotted in log scale, while the y-axis represents the predicted values, plotted in linear scale.
+Looking at the performance of the models by category, it is evident that the ML Objects-ML Captions Model is the most consistent and achieves the best performance across half of the categories, namely Drinking Water, Drying Clothes, Kitchens, and Washing Clothes.
+The ML Objects-Human Labels Model was the most effective for Front Doors.
+The Handwashing category was best handled by the ML Captions Model, while the Human Labels Model achieved the best results for Living Rooms and Places for Dinner.
+The observation that the ML Objects and ML Captions Model achieves highest performance aligns with the claim made by Zhang et al.
 (2021) on machines being better at selecting distinguishing features in the context of image segmentation.
@@
 Further experiments would need to be conducted to explore this.
-Furthermore, we compared the performance differences between the best and worst-performing models (overall), namely the ML Objects-ML Captions Model and the Human Model.
+Moreover, similar to the classification model, the all-annotations model worked best for African images, likely driven by the need for the annotations to encompass various features, including objects and actions.
+The images from Asia and the Americas were best handled by the Human Labels Model and the ML Captions-Human Labels Model, respectively.
+This shows how images from both regions have features indicative of income that are most likely not object-based, rather based on actions or nuanced descriptions of the environment.
+The fact that Asia required only Human Labels for its optimal performance suggests that the discriminative features for income still require full human intervention, and that although the ML annotations can still capture them, they do not perform as well, even when combined with the Human Labels.
+In contrast, the European images achieved best performance with the ML Objects-ML Captions Model.
+This suggests that the combined ML-based annotations are better at capturing the defining features and relationships to predict income for European images.
+This indicates a potential for bias in the ML annotations, potentially stemming from the training data of the ML models, which may be more representative of European content.
+However, another possibility is that the human annotators coincidentally managed to find better discriminating features for income prediction when annotating images from Asia.
+The best regression models for the action-based image categories are either the ML Captions Model or the ML Objects-ML Captions Model, indicating how the ML models are better at finding the defining features to predict income.
+It may be due to how the images contain more easily identifiable actions and objects that are indicative of income.
+On the other hand, image categories that are not action-based, such as Living Rooms and Places for Dinner are best handled by the Human Labels Model.
+This suggests that humans are better at identifying key features for income prediction when said features are likely to be based on nuanced descriptions.
+Furthermore, we compared the performance differences between the best and worst-performing models (overall), namely the ML Objects-ML Captions Model and the ML Captions-Human Model.
 The results of the Kolmogorov-Smirnov test showed that the difference is statistically significant ($p$ $<$ 0.001), suggesting that the ML Objects-ML Captions Model’s better performance is statistically reliable.
 Figure 5 illustrates a scatter plot of the regression model trained using the ML Objects and ML Captions to predict income on the test dataset.
-The data points are colour-coded by continent.
+Each data point represents an individual image sample from the dataset, which corresponds to a household.
+The points include associated attributes, such as region (continent), country, actual income, and predicted income.
+They are also colour-coded by continent.
 The blue dotted line represents the line of best fit and the red solid line represents the identity line.
-The plot reveals that most of the samples whose incomes are underestimated are mostly countries from Asia, with some samples from the Americas and Europe.
-On the other hand, most of the overestimated samples are from the Americas, with several samples from Asian and Europe.
-It is also important to note that none of the African samples deviate significantly from the identity line.
-The underestimation of samples from Asia may be attributed to the high number of samples within the lower to middle income range in the training set, resulting in the model skewing predictions for samples from Asia.
-Interestingly, when looking at the distribution of samples from the Americas, the income levels range from lower to middle income, as well, despite the overestimation by the model.
-This suggests that perhaps the features detected in the ML Objects and ML Captions for the Americas are not well-represented, causing inaccurate predictions.
-This is made worse due to the limited number of samples in the dataset, causing there to be less variety in income distribution for particular regions.
-Figure 4: Income Regression Model using ML Objects and ML Captions as Annotations Grouped by Continent
-Limitations.
-First, the experiments were done on only one dataset due to the limited availability of image datasets that contain features relating to socio-economic background, such as geographical region and income.
-Performing the experiments on another similar image dataset would make the results more reliable.
-Second, the number of image samples in the dataset that have been annotated by humans is rather small.
-Due to time constraints and limited resources, doing further data collection to obtain more human annotations was not feasible given the scope of this investigation.
-Having more annotated images would result in better generalisability of the models.
-Third, due to the small size of the annotated data, under-sampling was not performed for the regression task - unlike the classification model which had under-sampling embedded in the algorithm - which meant that there may be data imbalance when considering number of samples in each region.
-This could potentially be the reason as to why the performance of some regions is notably better than others.
-Lastly, there is a risk of misclassification and fault intolerance in this work, as the small dataset could affect the generalisability of the model and, therefore, the trends observed.
+The x-axis represents the ground truth of the income, which has been plotted in log scale for a clearer visualisation, whereas the y-axis represents the predicted income, plotted in its original value.
+The plot reveals that the samples whose incomes are underestimated come from all regions, with most of the Europe samples clustering towards the higher income area in the x-axis and the Africa samples clustering towards the lower income area in the x-axis.
+On the other hand, there are very few overestimated samples, most of which are only from Africa.
+The underestimation of samples may be attributed to the high number of samples within the lower to middle income range in the dataset.
+Interestingly, when looking at the distribution, samples from Africa tend to be closer to the red identity line, indicating a more accurate prediction, especially compared to Europe.
+This is expected as this is also reflected in the RMSE scores of the regression models in Table 4.
+The low variance of the RMSE values for African samples may be due to the features and relationships in the African samples captured in the ML Objects and ML Captions being more different from those in other regions, enabling the model to be more consistent in its predictions.
+On the other hand, the key features found in the annotations of the European samples may be more complex, resulting in more variability in predictions, as shown in the vertical spread of the European samples in the scatter plot, as well as the greater variance in the RMSE table.
+Limitations. Our study only considers certain ML models and certain types of images. We thus cannot claim that our observations would generalise beyond the data and model considered.
 Conclusion
@@
 We first look at RQ1: “How similar are human-generated and ML-generated annotations?”.
-It can be concluded that from a low-level perspective, i.e., what types of word occur, sentence structures, the ML Objects and ML Captions are similar to each other, but are relatively different to the human labels.
-This is expected due to underlying cognitive processes during the annotation stage.
-However, when looking at them from a high-level point of view, despite the apparent low similarity, they are alike in that they exhibit consistent similarity or dissimilarity regardless of the geographical region of the images, indicating lack of regional bias.
-This led to RQ2: “How do different combinations of annotations affect fairness in ML predictive models?” In the case of region classification, it was revealed that the human labels resulted in the best overall-performing model, as well as the most balanced performance across individual classes.
-It is additionally apparent that each region obtained best classification performance with models that utilised annotation combinations that included human labels.
-Similarly, for income regression, the model that has the best overall performance also has the most balanced performance across regions.
-However, in this case, the combination of ML Objects and ML Captions is what results in this enhanced performance.
-A similar trend continues in that the best performance for individual classes was obtained with models that contained ML Objects in its annotations.
-Considering these findings, it can be deduced that there is no particular combination of annotations that result in best overall and most balanced performance as that is highly dependent on the quality of the images and how well the human annotator or machine is able to select the best discriminative features.
-Additionally, due to imbalanced distributions of income across data samples and low number of samples for certain regions, the regression model is prone to underestimation or overestimation of income for certain regions.
-Moreover, our analysis suggests that it may be plausible that models with the best overall performance tend to also be more balanced on the class-level, and therefore, fairer.
-However, further studies need to be done to confirm this.
+It can be concluded that from a low-level perspective, i.e., what types of word occur, sentence structures, the ML Captions and Human Labels are similar to each other, but ML Objects are relatively different to either of them.
+This is expected due to similarities in underlying cognitive processes between ML models and human brains.
+However, when looking at them from a high-level point of view, despite the apparent low similarity between ML Objects and the other annotations, all of them are alike in that they exhibit consistent similarity or dissimilarity, even when considering the spread based on region, where all of them have the same trends, with Europe clustering together and showing a strong agreement across all annotation types.
+This led to RQ2: “How do different combinations of annotations affect fairness in ML predictive models?” In the case of region classification, it was revealed that different annotation combinations work best for different geographical regions and image categories.
+This points to the conclusions that all annotations are important, and machine-generated annotations cannot just replace human-generated ones.
+It is additionally apparent that each region obtained best classification performance with models that utilised annotation combinations that included ML Captions.
+In the case of income regression, the combination of ML Objects and ML Captions is what results in enhanced performance.
+Unlike the region classification, the best performance for individual classes was obtained from different annotation models.
+Considering these findings, it can be concluded that there is no particular combination of annotations that result in best region-based performance as that is highly dependent on the quality of the images and how well the human annotator or machine is able to select the best discriminative features.
 Implications.
@@
 First, despite their apparent differences, humans and machines may be more alike in their bias, or lack thereof, in the context of visual content.
-Their consistent similarity and dissimilarity of annotations across regions implies that their level of fairness is comparable.
+Their consistent similarity and dissimilarity of annotations across regions implies that their level of bias is comparable.
 This insight could steer researchers towards further advancing ML models to avoid as much bias as possible.
-Second, achieving a more balanced performance on the region level can help reduce biases in ML models.
-This would result in models that not only give accurate outcomes, but ones that are also fair.
-This would be highly beneficial in real-world scenarios where the model’s outcomes could disproportionately affect certain demographic groups.
-Third, due to machine annotations (ML Objects and ML Captions) performing better and fairer at the regression task, tasks in other domains could potentially benefit from using machine annotations as well.
+Second, due to machine annotations (ML Captions only and ML Objects-ML Captions) performing better overall at both the classification and regression tasks, tasks in other domains could potentially benefit from using machine annotations as well.
 However, this could result in a potential negative societal impact in that humans become overly reliant on machines, which may cause them to miss possible biases in the models that originated from the training data, miss errors due to assuming the superior performance of the ML models, or even cause difficulties in identifying accountability if the ML models’ prediction errors result in significant repercussions.
-This could also lead to potential misuse of the technology as it is possible that one intentionally uses a model that is potentially biased without attempting to address the fairness problems.
+This could also lead to potential misuse of the technology as it is possible that one intentionally uses a model that is potentially biased without attempting to address the underlying issues.
 Acknowledgements
+This work is partially supported by the Swiss National Science Foundation (SNSF) under contract number CRSII5_205975 and by an Australian Research Council (ARC) Future Fellowship Project (Grant No.
+FT240100022).
 ChatGPT and Claude AI were utilised to assist with brainstorming and to generate sections of this Work, including alternative phrasings and parts of code used for the experiments.

### v3

@@
 However, recently, language and multi-modal foundational models have been used to replace and scale-up human annotator’s efforts.
-This study compares human-generated and ML-generated annotations of images representing diverse socio-economic contexts.
+This study explores the similarity between human-generated and ML-generated annotations of images across diverse socio-economic contexts (RQ1) and their impact on ML model performance and bias (RQ2).
 We aim to understand differences in perception and identify potential biases in content interpretation.
 Our dataset comprises images of people from various geographical regions and income levels, covering various daily activities and home environments.
-We compare human and ML-generated annotations semantically and evaluate their impact on predictive models.
-Our results show highest similarity between ML captions and human labels from a low-level perspective, i.e., types of words that appear and sentence structures, but all three annotations are alike in how similar or dissimilar they perceive images across different regions.
-Additionally, ML Captions resulted in best overall region classification performance, while ML Objects and ML Captions performed best overall for income regression.
-The varying performance of annotation sets highlights the notion that all annotations are important, and that human-generated annotations are yet to be replaceable.
+ML captions and human labels show highest similarity at a low-level, i.e., types of words that appear and sentence structures, but all annotations are consistent in how they perceive images across regions.
+ML Captions resulted in best overall region classification performance, while ML Objects and ML Captions performed best overall for income regression.
+ML annotations worked best for action categories, while human input was more effective for non-action categories.
+These findings highlight the notion that both human and machine annotations are important, and that human-generated annotations are yet to be replaceable.
+Code and Datasetshttps://github.com/nardienapratama/Human-Machine-Perception-of-Visual-Content
 Introduction
@@
 In this study, we first investigate how similar annotations generated by humans and machines are, and explore possible justifications for their similarities or differences, in the context of how they think and perceive images.
-This leads to our first research question (RQ1): “How similar are annotations generated by humans compared to the annotations generated by ML models?” Moreover, due to the aforementioned biases that could appear in the annotations, we look into how different annotations affect unbiasedness in predictive models.
+This leads to our first research question (RQ1): “How similar are human-generated and ML-generated annotations?” Moreover, due to the aforementioned biases that could appear in the annotations, we look into how different annotations affect unbiasedness in predictive models.
 This leads to our second research question (RQ2): “How do different combinations of annotations affect the performance and bias in ML predictive models?” To do this, we use the annotations in vector representation to train predictive ML models.
@@
 Methodology
-Figure 1: Handwashing images annotated using the pre-trained Faster R-CNN model.
+Figure 1: Example images annotated with the pre-trained Faster R-CNN model. Left: Hand Washing. Right: Living Rooms.
 Experimental Set-Up
@@
 Dataset
-The dataset utilised in our study was obtained from Dollar Street, a project by Gapminder, and is available under a Creative Commons license CC BY 4.0 (Gapminder 2021).
-Dollar Street visualises global public data by arranging families based on their income levels, from the poorest on the left to the wealthiest on the right.
+The dataset utilised in this study comes from Dollar Street, a project by Gapminder, and is available under a Creative Commons license CC BY 4.0 (Gapminder 2021).
+Dollar Street visualises global public data by arranging families based on their income levels.
 It features photographs of everyday objects from more than 240 families in 50 countries, providing a representation of living conditions across different economic classes.
 The dataset contains 3,041 images, sourced from various geographical regions and income levels, consisting of the following image categories: Drinking Water, Drying Clothes, Front Doors, Hand Washing, Kitchens, Living Rooms, Places for Dinner, and Washing Clothes.
-These images include a diverse range of topics, covering both actions and static elements, such as home environments.
-This is to provide a good representation of different aspects of daily life, which is applicable across all cultures and income levels.
+The dataset captures a mix of actions and static elements, such as home environments, to reflect different aspects of daily life across cultures and income levels.
 Data Annotation
@@
 Data Preparation and Preprocessing
-Prior to obtaining all three annotation sets, the images were selected based on whether they successfully led to the generation of ML Objects with at least 50% confidence and ML Captions beforehand.
-Moreover, duplicates were removed and down-sampling was performed to ensure the images were balanced for all regions and income groups.
-The human annotations were processed to deal with inconsistencies.
-These steps involve noise/punctuation removal, conversion to lower case, auto-correcting misspellings, and segmenting strings that are concatenated.
+Prior to obtaining all three annotation sets, the images were selected based on whether ML Objects with at least 50% confidence and ML Captions could be generated.
+Duplicates were removed, and down-sampling was performed to ensure the images were balanced across regions and income groups.
+Human annotations were processed to address inconsistencies by removing noise/punctuation, converting to lower case, auto-correcting misspellings, and segmenting concatenated strings.
 The number of images in the dataset decreased at each step, starting at 3,041 images at the start, reducing to 2,797 after generating the ML Objects and Captions, then to 1,888 after removing duplicates and down-sampling, and finally to 1,886 unique samples after crowdsourcing.
 In total, 1,155 samples were discarded from the original dataset.
-We utilised a fine-tuned sentence transformer model to convert the three types of annotations into sentence embeddings.
-We opted to fine-tune an existing model instead of using an pre-trained model out of the box as we want the model and representations to be very focused to our data.
-We fine-tuned a pre-trained Sentence Transformer model, “all-MiniLM-L12-v2”, available at the Hugging Face Hub [^1] https://huggingface.co/sentence-transformers/all-MiniLM-L12-v2.
-The fine-tuning process was run over 5 epochs with a batch size of 16.
-The model was fine-tuned on the three annotation sets so that it can better capture the characteristics of the annotations, hence generating closer embeddings.
-The generated embeddings have a size of 384 dimensions, which is the default embedding size from the original pre-trained model.
-Exp. 1: Investigating Annotation Similarity (RQ1)
+We fine-tuned the pre-trained Sentence Transformer model, “all-MiniLM-L12-v2” (Reimers 2021), using code adapted from Hugging Face’s GitHub repository (Espejel 2022).
+The model was fine-tuned over 5 epochs with a batch size of 16 to better tailor its representations to our data, rather than using it off the shelf.
+Fine-tuning on the three annotation sets allowed the model to better capture their characteristics.
+The resulting sentence embeddings are 384-dimensional, consistent with the original pre-trained model.
+Experiment 1: Investigating Annotation Similarity (RQ1)
 To investigate the similarity between human, machine-generated object labels, and machine-generated captions, we utilised the embeddings that have been generated for the three types of annotations across all images.
@@
 These experiments are appropriate for addressing RQ1 as t-SNE, computation of similarity scores across pairs, and annotation length allow us to evaluate how similar the annotations are, whereas the pairwise scatter plot helps us derive trends and patterns between annotations, further evaluating their likeness to one another.
-Exp. 2: Exploring How Annotations Impact Performance and Bias in Predictive Models (RQ2)
-This section emphasises the diversity of model prediction outcomes over overall model effectiveness.
-These experiments aim to evaluate the distribution of prediction errors across different regions and categories, allowing us to identify potential discrimination, e.g., higher prediction errors for certain subsets of the population as compared to the error on others.
+Experiment 2: Exploring How Annotations Impact Performance and Bias in Predictive Models (RQ2)
+This section emphasises the diversity of model prediction outcomes over overall performance.
+These experiments evaluate the distribution of prediction errors across different regions and categories, allowing us to identify potential discrimination, e.g., higher prediction errors for certain subsets of the population as compared to the error on others.
 To investigate how the three annotation sets affect predictive models, we use the vector embeddings of every annotation as training data for two supervised learning models.
@@
 The COCO dataset, which the ML object detection model is trained on, does not extend to action recognition, which is not relevant due to the object-oriented images in the dataset, thereby inherently limiting the ML-generated object annotations to object identification.
-The ML-generated captions would be more descriptive than the object labels and although they may not be as specific as the human annotations, they are more likely to contain actions and objects, which may also be present in Human Labels.
+Moreover, the ML-generated captions would be more descriptive than the object labels and although they may not be as specific as the human annotations, they are more likely to contain actions and objects, which may also be present in Human Labels, hence the clustering of the ML-generated captions and the Human Labels.
 Additionally, there are apparent syntactic differences between the annotations - the ML Objects consists of lists of objects per image; the ML Captions consist of a phrase per image; while the Human Labels are a list of words and phrases.
@@
 It is important to note that when comparing ML Objects to any of the other annotations, the similarity scores are relatively low.
-To further investigate length statistics per image, Human Labels averaged 7.18 words per worker, ML Captions averaged 8.62 words per caption, and ML Objects averaged 4.36 object labels per prediction.
-The length of ML Objects is determined by the number of objects in the predictions rather than the average number of words as ML Objects consists of a discrete list of items.
+To further investigate length statistics per image, Human Labels averaged 7.18 words per worker, ML Captions averaged 8.62 words per caption, and ML Objects averaged 4.36 object labels per inference run.
+The length statistic for ML Objects is determined by the number of object labels rather than the number of words because each predicted label corresponds to a single object.
 Figure 4: Pairwise relationships between similarity scores.
@@
 The plot illustrates how similarity scores between two annotation categories correlate to another pair’s scores for a given image.
-The plot is colour-coded by region to allow further analysis of similarity scores by geographical location.
+The plot is colour-coded by region to allow further analysis of pairwise similarity scores by geographical location.
 The analysis reveals that most regions show no particular trends or patterns, with one exception: images from Europe tend to cluster in the higher similarity range (top right area) of the plots, indicating stronger agreement between all annotation types for European images.
@@
 This aligns with the previous observation that ML Captions and Human Labels have the highest average similarity score (0.69).
-It can also be seen that all three pairs show the same amount of spread of scores across all subplots, which correspond to the variance values of the mean similarity scores between each pair (0.12).
+It can also be seen that all three pairs show the same amount of spread of scores across all subplots, which correspond to the standard deviation values of the mean similarity scores between each pair (0.12).
 The lack of discernible patterns in the way the scores are distributed across the Americas, Africa, and Asia suggests that images from these regions have little to no effect on how similar two sets of annotations will be.
@@
 We hypothesise that this due to the fact that both humans and machine-based models are more familiar with Western societies images based on their past experience or training.
-Moreover, the fact that the ML Objects annotations have a relatively low similarity score indicates that the ML object detection model may be detecting objects that do not appear in the corresponding ML Captions or Human Labels, or failing to identify objects that do.
+Moreover, the fact that the ML Objects annotations have a relatively low similarity score with the other annotations indicates that the ML object detection model may be detecting objects that do not appear in the corresponding ML Captions or Human Labels, or failing to identify objects that do.
 This reasoning is consistent with the findings of Zhang et al.
@@
 The variation in optimal models across the image categories is understandable.
-Images from each of the categories may contain varying objects and/or actions that are associated with certain geographical regions, e.g., for theHand Washing category, images of buckets may appear more frequently in samples from Africa, whereas images of sinks may appear more in samples from Europe.
+Images from each of the categories may contain varying objects and/or actions that are associated with certain geographical regions, e.g., for Hand Washing, images of buckets may appear more frequently in samples from Africa, whereas images of sinks may appear more in samples from Europe.
 This explains why certain annotations are better than others at capturing relationships between the objects or actions and the regions.
-The best-performing models for five of the categories (Drying Clothes, Kitchens, Living Rooms, Places for Dinner, and Washing Clothes) are the ones with a combination of at least two annotation sets.
+The best-performing models for five out of eight categories (Drying Clothes, Kitchens, Living Rooms, Places for Dinner, and Washing Clothes) are the ones with a combination of at least two annotation sets.
 This may be due to how the images from these categories may have discriminative features that cannot be captured with just one annotation set.
-Using the Welch’s t-test, we found that there are statistically significant differences between the overall F1-scores from 10 independent runs across the ML Captions Model and the Human Labels Model at $p$-value $<$ $1.15\times 10^{-7}$.
+Using the Welch’s t-test, we found that there are statistically significant differences between the overall F1-scores from 10 independent runs across the ML Captions Model (overall best-performing) and the Human Labels Model (overall worst-performing) at $p$-value $<$ $1.15\times 10^{-7}$.
 This statistically significant value suggests that the ML Captions Model’s better performance is statistically reliable.
+Figure 5: Income Regression Model using ML Objects and Captions as Annotations Grouped by Continent.
+The x-axis represents the ground truth values, plotted in log scale, while the y-axis represents the predicted values, plotted in linear scale.
 Table 4 shows the Root Mean Square Error (RMSE) for the AdaBoost Regression algorithm used to predict income grouped by region, while Table 5 shows the scores grouped by image category.
@@
 The best-performing models differ across all regions.
-The model containing all annotations was the most effective for African images (561.42), similar to the classification model, while the Asian images achieved the best results with the Human Labels Model (1408).
+The model containing all annotations was the most effective for African images (561.42), similar to the classification model, while the Asian images achieved the best results with the Human Labels Model (1408.57).
 Europe saw optimal performance from the ML Objects and ML Captions Model (2777.27), and the Americas was best handled by the ML Captions and Human Labels Model (1578.29).
 Moreover, it is worth highlighting that Asia is the only region where a stand-alone annotation set achieved the best performance.
-Figure 5: Income Regression Model using ML Objects and Captions as Annotations Grouped by Continent.
-The x-axis represents the ground truth values, plotted in log scale, while the y-axis represents the predicted values, plotted in linear scale.
-Looking at the performance of the models by category, it is evident that the ML Objects-ML Captions Model is the most consistent and achieves the best performance across half of the categories, namely Drinking Water, Drying Clothes, Kitchens, and Washing Clothes.
+Looking at the performance of the models by category, the ML Objects-ML Captions Model is the most consistent and achieves the best performance across half of the categories, namely Drinking Water, Drying Clothes, Kitchens, and Washing Clothes.
 The ML Objects-Human Labels Model was the most effective for Front Doors.
@@
 (2021) on machines being better at selecting distinguishing features in the context of image segmentation.
-In our case of income regression, this also applies, although only for ML Objects rather than ML Captions.
-This could be attributed to certain objects in the images being associated with certain income levels, therefore improving the model’s ability to predict income.
+In this case of income regression, the superior performance of the ML-based annotations could be attributed to certain objects in the images, as well as descriptive terms indicating the state of the environment (e.g., clean, dirty, old), being associated with certain income levels, therefore improving the model’s ability to predict income.
 Further experiments would need to be conducted to explore this.
@@
 This suggests that the combined ML-based annotations are better at capturing the defining features and relationships to predict income for European images.
-This indicates a potential for bias in the ML annotations, potentially stemming from the training data of the ML models, which may be more representative of European content.
-However, another possibility is that the human annotators coincidentally managed to find better discriminating features for income prediction when annotating images from Asia.
-The best regression models for the action-based image categories are either the ML Captions Model or the ML Objects-ML Captions Model, indicating how the ML models are better at finding the defining features to predict income.
+This suggests potential bias in the ML annotations, possible due to Western-centric training data.
+Alternatively, human annotators may have found better discriminative features for income prediction in images from Asia.
+The best regression models for the action-based image categories, e.g., Drying Clothes, Hand Washing, are ones with ML-based annotations, indicating how the ML annotations are better at finding the defining features to predict income.
 It may be due to how the images contain more easily identifiable actions and objects that are indicative of income.
-On the other hand, image categories that are not action-based, such as Living Rooms and Places for Dinner are best handled by the Human Labels Model.
+On the other hand, most image categories that are not action-based, such as Living Rooms and Places for Dinner, are best handled by models containing human annotations.
 This suggests that humans are better at identifying key features for income prediction when said features are likely to be based on nuanced descriptions.
@@
 The results of the Kolmogorov-Smirnov test showed that the difference is statistically significant ($p$ $<$ 0.001), suggesting that the ML Objects-ML Captions Model’s better performance is statistically reliable.
-Figure 5 illustrates a scatter plot of the regression model trained using the ML Objects and ML Captions to predict income on the test dataset.
+Figure 5 illustrates a scatter plot of the income regression model trained using the ML Objects and ML Captions (best overall) on the test dataset.
 Each data point represents an individual image sample from the dataset, which corresponds to a household.
@@
 The blue dotted line represents the line of best fit and the red solid line represents the identity line.
-The x-axis represents the ground truth of the income, which has been plotted in log scale for a clearer visualisation, whereas the y-axis represents the predicted income, plotted in its original value.
-The plot reveals that the samples whose incomes are underestimated come from all regions, with most of the Europe samples clustering towards the higher income area in the x-axis and the Africa samples clustering towards the lower income area in the x-axis.
-On the other hand, there are very few overestimated samples, most of which are only from Africa.
-The underestimation of samples may be attributed to the high number of samples within the lower to middle income range in the dataset.
-Interestingly, when looking at the distribution, samples from Africa tend to be closer to the red identity line, indicating a more accurate prediction, especially compared to Europe.
-This is expected as this is also reflected in the RMSE scores of the regression models in Table 4.
-The low variance of the RMSE values for African samples may be due to the features and relationships in the African samples captured in the ML Objects and ML Captions being more different from those in other regions, enabling the model to be more consistent in its predictions.
+The x-axis represents the ground truth of the income, which has been plotted in log scale for clearer visualisation, whereas the y-axis represents the predicted income, plotted in its original value.
+The plot reveals that the incomes of most samples were overestimated, with most of the Africa samples clustering towards the lower income area in the x-axis.
+On the other hand, the three other regions are more balanced, although there is a great portion of underestimated samples, most of which are from Europe.
+Although the data was sampled evenly across continents and balanced across income groups within each continent, the ground truth income levels of most non-African samples are higher than the African samples’.
+This is evident in the scatter plot, where non-African samples are predominantly positioned to the right of African points.
+This means that the overall average income across all samples is higher than most African samples, which may attribute to their overestimation.
+The low variance in the predicted income of African samples may be due to the features in these samples, captured in the ML Objects and ML Captions, being simpler to capture and understand during model training.
+This simplicity enables the model to be more consistent in its predictions.
 On the other hand, the key features found in the annotations of the European samples may be more complex, resulting in more variability in predictions, as shown in the vertical spread of the European samples in the scatter plot, as well as the greater variance in the RMSE table.
@@
 We first look at RQ1: “How similar are human-generated and ML-generated annotations?”.
-It can be concluded that from a low-level perspective, i.e., what types of word occur, sentence structures, the ML Captions and Human Labels are similar to each other, but ML Objects are relatively different to either of them.
+From a low-level perspective, i.e., what types of word occur and sentence structures, the ML Captions and Human Labels are similar to each other, but ML Objects are relatively different to either of them.
 This is expected due to similarities in underlying cognitive processes between ML models and human brains.
-However, when looking at them from a high-level point of view, despite the apparent low similarity between ML Objects and the other annotations, all of them are alike in that they exhibit consistent similarity or dissimilarity, even when considering the spread based on region, where all of them have the same trends, with Europe clustering together and showing a strong agreement across all annotation types.
-This led to RQ2: “How do different combinations of annotations affect fairness in ML predictive models?” In the case of region classification, it was revealed that different annotation combinations work best for different geographical regions and image categories.
-This points to the conclusions that all annotations are important, and machine-generated annotations cannot just replace human-generated ones.
-It is additionally apparent that each region obtained best classification performance with models that utilised annotation combinations that included ML Captions.
-In the case of income regression, the combination of ML Objects and ML Captions is what results in enhanced performance.
-Unlike the region classification, the best performance for individual classes was obtained from different annotation models.
-Considering these findings, it can be concluded that there is no particular combination of annotations that result in best region-based performance as that is highly dependent on the quality of the images and how well the human annotator or machine is able to select the best discriminative features.
+However, from a high-level point of view, despite the apparent low similarity between ML Objects and the other annotations, all exhibit consistent trends, even when considering the spread based on region, with Europe clustering together and showing a strong agreement across all annotation types.
+This led to RQ2: “How do different combinations of annotations affect the performance and bias in ML predictive models?” For region classification, different annotation combinations work best for various regions and image categories.
+Models that included ML Captions achieved the best performance in each region.
+Most image categories performed best with annotations containing at least two types.
+For income regression, using ML Object Labels and ML Captions enhanced performance, particularly across many image categories.
+The best performance for each region was obtained from a wide variety of annotations.
+Similar to region classification, performance was highest in Africa when all annotation types were used, likely due to the images containing object/action cues and nuanced environmental context.
+ML-based annotations worked well for action image categories, while human input benefited non-action ones.
+Using diverse annotations led to more consistent predictions, as seen in Africa’s lower variance in the scatter plot.
+Considering these findings, it can be concluded that all annotations are important and machine-generated annotations cannot fully replace human-generated ones.
 Implications.
 These findings have several implications.
-First, despite their apparent differences, humans and machines may be more alike in their bias, or lack thereof, in the context of visual content.
-Their consistent similarity and dissimilarity of annotations across regions implies that their level of bias is comparable.
-This insight could steer researchers towards further advancing ML models to avoid as much bias as possible.
-Second, due to machine annotations (ML Captions only and ML Objects-ML Captions) performing better overall at both the classification and regression tasks, tasks in other domains could potentially benefit from using machine annotations as well.
-However, this could result in a potential negative societal impact in that humans become overly reliant on machines, which may cause them to miss possible biases in the models that originated from the training data, miss errors due to assuming the superior performance of the ML models, or even cause difficulties in identifying accountability if the ML models’ prediction errors result in significant repercussions.
-This could also lead to potential misuse of the technology as it is possible that one intentionally uses a model that is potentially biased without attempting to address the underlying issues.
+First, despite their apparent differences, humans and machines may share similar levels of bias in visual content.
+Their consistent patterns across regions show comparable tendencies.
+This points to the need for further advancement to reduce bias in ML models.
+Second, since machine annotations (ML Captions only and ML Objects-ML Captions) perform better overall at both the classification and regression tasks respectively, they may be beneficial in other domains.
+However, being overly reliant on machine annotations may have negative societal impacts, such as missing possible biases in the models that originated from the training data, assuming ML models are always superior, or struggling to identify accountability if errors lead to significant consequences.
+This could also lead to potential misuse of the technology, where biased models are intentionally used without addressing underlying issues.
 Acknowledgements